### v1

# Consistent Validation for Predictive Methods in Spatial Settings

Abstract
Spatial prediction tasks are key to weather forecasting, studying air pollution, and other scientific endeavors. Determining how much to trust predictions made by statistical or physical methods is essential for the credibility of scientific conclusions. Unfortunately, classical approaches for validation fail to handle mismatch between locations available for validation and (test) locations where we want to make predictions. This mismatch is often not an instance of covariate shift (as commonly formalized) because the validation and test locations are fixed (e.g., on a grid or at select points) rather than i.i.d. from two distributions. In the present work, we formalize a check on validation methods: that they become arbitrarily accurate as validation data becomes arbitrarily dense. We show that classical and covariate-shift methods can fail this check. We instead propose a method that builds from existing ideas in the covariate-shift literature, but adapts them to the validation data at hand. We prove that our proposal passes our check. And we demonstrate its advantages empirically on simulated and real data.
1 Introduction
Researchers are often interested in making predictions in a spatial setting. For instance, scientists would like to predict sea surface temperature (SST) for weather forecasting and climate research (Minnett, 2010), predict air pollution at population centers to better understand the effect of pollution on health outcomes such as kidney disease (Remigio et al., 2022), or predict the prevalence of an invasive species for ecological management (Barbet-Massin et al., 2018). Characterizing the reliability of these predictions is key to understanding their suitability for certain downstream applications; e.g., Minnett (2010) describe acceptable SST error tolerances for weather forecasting. Estimating prediction accuracy can also be used to decide among a collection of predictive methods, as in Shabani et al. (2016).
Importantly, in the spatial setting, predictive methods need not always arise from a statistical or machine learning approach built using training data. The predictive method is often a complex physical model provided by a third party (Remigio et al., 2022; Minnett, 2010; Gupta et al., 2018). Or it could combine physics and data-driven models (Banzon et al., 2016; Werner et al., 2019; Özkaynak et al., 2013).
In any of these cases, it is common to estimate the performance of a predictive method by using a set of validation data. More precisely, we are ultimately interested in predicting a response at what we will call test sites; in the SST example above (Minnett, 2010), the test sites are points on a grid (often called a map), or in the air pollution example (Remigio et al., 2022), the test sites are 28 counties in the US Northeast. We can make predictions at the test data, but we do not have access to direct observations of the responses there. We do have observed responses in the validation data, which we assume were not used in forming the predictive method being evaluated; in the SST example, scientists have SST observations taken by boats and buoys as validation data. So our aim is to estimate the average loss (i.e. risk) at the testing data using the validation data. We will see that many popular or natural approaches fail at this task.
One widely used approach, called the holdout, [^1] The name originally referred to “holding out” data for validation, with the remainder of available data going toward training. While we maintain the naming convention, we emphasize that in the present development there need not be any training data. Minnett (2010); Gupta et al. (2018); Duan et al. (2019), among many others, use this approach. estimates the test risk by taking the empirical average of the validation loss. When the validation and test data are independent and identically distributed (i.i.d.) from the same distribution, the holdout has a rigorous justification (Devroye, 1976; Langford, 2005). But in spatial problems, the validation and test sites need not be similarly dispersed, all data may be spatially correlated, and the test sites are often fixed rather than random; recall the grid or point prediction examples above. Indeed, Roberts et al. (2017) have observed problems with the holdout in practice. In the special case where the predictive method is data-driven, some authors (e.g. Telford & Birks, 2005) have suggested choosing holdout validation sites far from the training data sites. But we expect this proposal to still suffer from the problems just mentioned, and indeed simulation studies suggest it can misestimate test risk (Ploton et al., 2020; De Bruin et al., 2022).
Another natural idea is to use covariate-shift approaches to handle potential mismatch between validation and test sites (Sarafian et al., 2020; De Bruin et al., 2022). However, the covariate-shift literature generally assumes validation sites are drawn i.i.d. from one distribution, test sites are i.i.d. from another, and the density ratio between these two distributions exists and is bounded. For both the grid and point examples, these last two assumptions are inappropriate.
In what follows, we start by laying out a precise formulation of the prediction validation task in the spatial setting (Section 2). We formalize a desirable property for test-risk estimators: that, if arbitrarily dense validation data accrues in a region including the test points, the test-risk estimate should become arbitrarily accurate (Section 3). We prove that both the holdout estimator and an estimator advocated in the covariate-shift literature (Loog, 2012; Portier et al., 2023) fail to satisfy this spatial consistency property (Section 4). We propose instead to build on the $k$-nearest neighbor estimator of Loog (2012). In particular, Loog (2012); Portier et al. (2023) advocated fixing $k=1$ for covariate-shift problems. We instead derive an upper bound on the error of the general-$k$ estimator for estimating test risk (Section 5.1); crucially, our bound is conditional on the test and validation sites. We prove that choosing $k$ by optimizing our upper bound yields an estimator with our desired spatial consistency property (Section 5.2). Unlike covariate-shift results (e.g. Portier et al., 2023), our results are directly applicable to problems where the test sites are most reasonably thought of as fixed. We illustrate the accuracy and practicality of our proposed method in simulated and real data analyses (Section 6), including tasks in both grid and point prediction. We discuss further related work in App. A.
2 Estimating Test Risk in a Spatial Problem
We now formalize risk predictions at test points in a spatial setting. We assume each data point occurs at a spatial location $S\in\mathcal{S}$, where the spatial domain $(\mathcal{S},d_{\mathcal{S}})$ is a metric space. Each data point has observed covariates $X\in\mathcal{X}$ and a response $Y\in\mathcal{Y}$. We treat the covariates as a fixed spatial field, $\chi:\mathcal{S}\to\mathcal{X}$; i.e., the covariates at a point are exactly specified by applying $\chi$ at the spatial location of the point.
2.1 Test risk of a spatial predictive method
We assume we have access to a predictive method $h:(\mathcal{S},\mathcal{X})\to\mathcal{Y}$. Since we assume covariates are determined by their spatial location, we will find it convenient to define $h^{\chi}:\mathcal{S}\to\mathcal{Y}$ to be the prediction made by $h$ at the location $S$: $h^{\chi}(S)=h(S,\chi(S))$. We suppose that practitioners would like to use $h$ to predict the response at a set of test sites where the response is currently unknown. We will find it useful to collect the $M^{\textup{{test}}}$ test data points, including true but unobserved responses, in $D^{\textup{{test}}}=(S^{\textup{{test}}}_{m},X^{\textup{{test}}}_{m},Y^{\textup{{test}}}_{m})_{m=1}^{M^{\textup{{test}}}}$.
To quantify performance of a predictive method, we need a loss. We will assume the loss is bounded, as is typically the case for practically-bounded responses; cf. temperature, pressure, or other physical quantities. [^2] Loss is also bounded for classification error or robust regression cases such as trimmed squared loss or Tukey’s biweight loss.
Assumption 2.1 ($\Delta$-bounded Loss).
The loss $\ell$ is a non-negative, $\Delta$-bounded function. That is $\ell:\mathcal{Y}\times\mathcal{Y}\to[0,\Delta]$.
Due to practical considerations such as measurement error, the response at a test point is usefully modeled as random. To summarize loss over this randomness, it is standard to consider expected loss (a.k.a. risk) at the testing data. To define this expectation, we need to make assumptions about the data-generating process. It is typical in the non-spatial setting to assume responses are i.i.d. conditional on covariates. In the spatial setting, the i.i.d. assumption is inappropriate since it ignores spatial location. We instead assume that the response variable may be a function of the spatial location it is observed at, the covariates at that location, and i.i.d noise:
Assumption 2.2 (Data Generating Process: Test Data).
Let $j=\text{test}$. Let $\chi:\mathcal{S}\to\mathcal{X}$ be a fixed function. For $1\leq m\leq M^{j}$, $X^{j}_{m}=\chi(S^{j}_{m})$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m},\epsilon_{m}^{j})$ with $f:\mathcal{S}\times\mathcal{X}\to\mathcal{Y}$ and $\epsilon_{m}^{j}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P_{\epsilon}$ real-valued random variables.
A widely studied special case of Assumption 2.2 considers additive, homoskedastic noise: namely, $\mathcal{Y}\subset\mathbb{R}$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+\epsilon_{m}^{j}$. Assumption 2.2 is more general; for example, it allows the noise to be scaled by a continuous, deterministic function of $\mathcal{S}$: $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+g(S^{j}_{m})\epsilon_{m}^{j}$.
[…]
$Q^{\textup{{test}}}\coloneqq(1/M^{\textup{{test}}})\sum\nolimits_{m=1}^{M^{\textup{{test}}}}\delta_{S^{\textup{{test}}}_{m}},$ (1)
where for $S\in\mathcal{S}$, $\delta_{S}$ denotes a Dirac measure at $S$. In this case, we call the resulting risk, written out in Def. 2.3 next, the test risk of $h$.
Definition 2.3.
Give a collection of test points $(S^{\textup{{test}}}_{m})_{m=1}^{M^{\textup{{test}}}}$ and a predictive method $h$, we define the test risk of $h$ by
$R^{\chi}_{Q^{\textup{{test}}}}(h)=(1/M^{\textup{{test}}})\sum\nolimits_{m=1}^{M^{\textup{{test}}}}e^{\chi}_{h}(S^{\textup{{test}}}_{m}).$ (2)
2.2 Estimating test risk
To estimate test risk, we assume we have access to $N^{\textup{{val}}}$ validation data points, collected in $D^{\textup{{val}}}=(S^{\textup{{val}}}_{n},X^{\textup{{val}}}_{n},Y^{\textup{{val}}}_{n})_{n=1}^{N^{\textup{{val}}}}$. We assume practitioners did not make use of either the validation or test data when constructing the predictive method.
As an example, the common holdout estimator of test risk uses the empirical average of the validation risk:
$\hat{R}^{\chi}_{\textup{{Hold}}}(h)\coloneqq(1/N^{\textup{{val}}})\sum\nolimits_{n=1}^{N^{\textup{{val}}}}\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n})).$ (3)
For validation data to provide information about test risk, we need regularity assumptions. First, we make a standard assumption that validation data follows the same data-generating process as test data. And second, we assume a form of smoothness across the spatial locations.
Assumption 2.4 (Data Generating Process: Validation Data).
Assumption 2.2 remains true when we take $j=\text{val}$, with the same $f,\chi$, and $P_{\epsilon}$ as for $j=\text{test}$.
Assumption 2.5 ($L$-Lipschitz).
There exists $L>0$ such that $\forall S,S^{\prime}\in\mathcal{S}$, $|e^{\chi}_{h}(S)-e^{\chi}_{h}(S^{\prime})|\leq Ld_{\mathcal{S}}(S,S^{\prime})$.
Assumption 2.5 arises naturally in many spatial tasks. For example, consider $\mathcal{Y}\subset[0,1]$, $\mathcal{S}=(\mathbb{R}^{d},\|\cdot\|_{2})$, and squared loss. Suppose we are in the homoskedastic, additive noise setting. Suppose $\mathbb{E}[f(S,\chi(S))+\epsilon|S]$ is $L_{Y}$-Lipschitz and $h^{\chi}$ is $L_{h}$-Lipschitz. Then Assumption 2.5 holds with $L=2(L_{Y}+L_{h})$; see Prop. B.1.
3 We Want Consistent Estimators
Once we have an estimator of test risk, it remains to check if that estimator performs well. We next formalize one natural check on performance: namely, estimators should become arbitrarily accurate if given validation data that is arbitrarily dense in the spatial domain. This check is analogous to traditional consistency in the i.i.d. data setting.
To that end, recall the fill distance as a measure of discrepancy between two sets of points. [^3] The fill distance is not a distance in the mathematical sense since it is asymmetric and can equal 00 in cases when its two arguments are not exactly equal.
Definition 3.1 (Fill distance, Cressie 2015, §5.8, Wendland 2004, Definition 1.4).
Let $(\mathcal{S},d_{\mathcal{S}})$ be a metric space and $\Psi_{1},\Psi_{2}\subset\mathcal{S}$. The fill distance of $\Psi_{1}$ in $\Psi_{2}$ is,
$\zeta(\Psi_{1};\Psi_{2}):=\sup\nolimits_{S_{2}\in\Psi_{2}}\inf\nolimits_{S_{1}\in\Psi_{1}}d_{\mathcal{S}}(S_{1},S_{2}).$ (4)
That is, the fill distance of $\Psi_{1}$ in $\Psi_{2}$ reflects the maximum distance from a point in $\Psi_{2}$ to the nearest point in $\Psi_{1}$. [^4] The fill distance is often defined with $\Psi_{2}=\mathcal{S}$. We will sometimes require the more general definition.
In the spatial statistics literature, infill asymptotics describes cases where data are gathered over a compact spatial domain in such a way that the fill distance of the data to its domain tends to 00 (Cressie, 2015, Chapter 5.8). We will say an estimator is consistent for the test risk under infill asymptotics if – for any $Q^{\textup{{test}}}$, $\chi$, and $h$ satisfying our assumptions above – the estimator converges in probability to $R^{\chi}_{Q^{\textup{{test}}}}(h)$.
Definition 3.2 (Consistency of Test Risk Estimation Under Infill Asymptotics).
Fix a predictive method $h$ and a test measure $Q^{\textup{{test}}}$. Take Assumptions 2.2, 2.1, 2.4 and 2.5. Consider an infinite sequence of validation sets of increasing size: $(D^{\textup{{val}}}_{N})_{N=1}^{\infty},D^{\textup{{val}}}_{N}=(S^{\textup{{val}}}_{N},X^{\textup{{val}}}_{N},Y^{\textup{{val}}}_{N})_{n=1}^{N}$ such that when $N^{\prime}<N$, the first $N^{\prime}$ points of $D^{\textup{{val}}}_{N}$ are $D^{\textup{{val}}}_{N^{\prime}}$. Suppose $\lim_{N\to\infty}\zeta(S^{\textup{{val}}}_{1:N},\mathcal{S})=0$. Let $\hat{R}_{N}$ be an estimator constructed from the validation data $D^{\textup{{val}}}_{N}$. We say that the estimator $\hat{R}_{N}$ is consistent for the test risk under infill asymptotics if $\lim_{N\to\infty}\mathrm{Pr}(|\hat{R}_{N}-R^{\chi}_{Q^{\textup{{test}}}}(h)|\geq\epsilon)=0.$
That is, as validation data fills the spatial domain, the estimator should converge to the test risk – no matter the composition of test sites. Note that we do not require convergence to be uniform over test measures. Also note that our assumption that fill distance tends to zero is generally weaker than an assumption that the validation sites are drawn i.i.d. from a distribution with Lebesgue density supported on the spatial domain. Reznikov & Saff (2015, Theorem 2.1) showed an implication relationship between these assumptions in a much more general setting for much more general metric spaces, and Vacher et al. (2021, Lemma 12) discusses the special case for the unit cube. Next, we present a finite sample version of this implication, with an advantage relative to past work that we keep track of all constants.
Proposition 3.3 (Independent and Identically Distributed Data Satisfies an Infill Assumption).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$, and $P$ has Lebesgue density lower bounded by $c>0$ over $[0,1]^{d}$. Let $B_{d}=\pi^{d/2}/\Gamma(\frac{d}{2}+1)$ be the volume of the $d$-dimensional Euclidean unit ball. For any $\delta\in(0,1)$ there exists an $n_{0}$ such that for all $N^{\textup{{val}}}\geq n_{0}$ with probability at least $1-\delta$
$\zeta(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},[0,1]^{d})\leq\left(\tfrac{4^{d}}{cN^{\textup{{val}}}B_{d}}\left(\log\tfrac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}\right)\right)^{1/d}.$ (5)
We prove Prop. 3.3 in Section B.2. The right-hand side of this bound is $O((\log N^{\textup{{val}}}/N^{\textup{{val}}})^{1/d})$, and so the fill distance converges to zero in probability under these assumptions.
Consistency under infill asymptotics is a minimal desirable property.
Much like traditional consistency, we emphasize that consistency as described here is just a single check among many. For instance, often practitioners will be interested in extrapolation far from observed data, which is not modelled by infill asymptotics and will need to be considered separately. Our only supposition here is that we will generally prefer test-risk estimators that satisfy consistency under infill asymptotics to those that do not.
4 Current Estimators Exhibit Inconsistency
Even though consistency under infill asymptotics is a minimal property, we next prove that principle existing test-risk estimators fail to satisfy it.
4.1 Inconsistency of the Holdout
The holdout estimator of test risk is well-justified and consistent in the traditional case of i.i.d. test and validation data (Devroye, 1976; Langford, 2005). But we next show that it need not be consistent for spatial problems.
Proposition 4.1 (Counterexample to the consistency of holdout).
Consider a spatial location $S$ and a single test point at this location $(S,X,Y)$, with covariate $X$ as at the start of Section 2 and response $Y$ as in Assumption 2.2. Suppose the test risk at such a point (Def. 2.3 with $M^{\textup{{test}}}=1$) is not constant as a function of $S$. Then there exists a test set (of size one) such that $\hat{R}^{\chi}_{\textup{{Hold}}}$ is not a consistent estimator of the test risk on that test set.
Intuition for Inconsistency of the Holdout.
Prop. 4.1 relies on the key observation that $\hat{R}^{\chi}_{\textup{{Hold}}}$ has no dependence on the test task. So unless all test tasks have the same risk (which will be true only in unusually simplistic settings), it cannot estimate them all consistently. One might think of the holdout estimator as generally exhibiting non-trivial bias since it averages loss across the validation sites when we really care about loss at the test sites. We give a formal proof following this intuition in Section B.3.1.
4.2 Inconsistency of 1-Nearest Neighbor
Because of the assumed regularity in the error function (Assumption 2.5), it is natural to estimate the error at a test site using nearby validation points. Loog (2012) proposed risk estimators using $k$-nearest neighbors in the context of covariate shift. Both Loog (2012) and Portier et al. (2023) advocated for the use of 1-nearest neighbor (1NN) in the covariate-shift setting, with the latter providing theoretical justifications under standard covariate-shift assumptions. However, we next produce common spatial examples where the 1NN estimator exhibits inconsistency.
Nearest Neighbor Estimator.
We start by reviewing a general $k$-nearest neighbor estimator, which we revisit later. Define the $k$-nearest neighbor radius of a point $S\in\mathcal{S}$ as $\tau^{k}(S)=\inf\{a\in\mathbb{R}:|S^{\textup{{val}}}_{1:N^{\textup{{val}}}}\cap B(S,a)|\geq k\},$ where $B(S,a)$ denotes the ball of radius $a$ centered at $S$. Then the $k$-nearest neighbor set [^5] We will state our results for the version of nearest neighbors where ties are resolved by including all equidistant points. However, our analysis holds for arbitrary tie-breaking methods. of a point $S\in\mathcal{S}$ is
$A^{k}(S)=\{1\leq n\leq N^{\textup{{val}}}:\,S^{\textup{{val}}}_{n}\in B(S,\tau^{k}(S))\}.\,$ (6)
As long as $1\leq k\leq N^{\textup{{val}}}$, $A^{k}(S)$ contains at least $k$ points: the $k$ nearest neighbors to $S$ in the validation set. It may be larger than $k$ if multiple points are equidistant from $S$.
Definition 4.2.
The $k$-nearest neighbor test-risk estimator is defined by
$\hat{R}^{\chi}_{w^{\textup{{NN}},k}}$ $(h)=\sum\nolimits_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n})),\textrm{ where}$ (7)
$\!w^{\textup{{NN}},k}_{n}$ $\!=\!\tfrac{1}{M^{\textup{{test}}}}\sum\nolimits_{m=1}^{M^{\textup{{test}}}}\tfrac{1}{|A^{k}(S^{\textup{{test}}}_{m})|}\mathrm{1}\{S^{\textup{{val}}}_{n}\!\in\!A^{k}(S^{\textup{{test}}}_{m})\}.$ (8)
Loog (2012) proposed weighting the loss function in this way when training a model under covariate shift. Portier et al. (2023) analyzed a similar approach, in which validation points are sampled with probabilities corresponding to the weights in Eqn. 8, for mean estimation. Portier et al. (2023) made standard covariate shift assumptions of i.i.d. validation sites, i.i.d. test sites, and a bounded density ratio between the validation and test distributions.
Our inconsistency result (Prop. 4.3 below) relies on the observation that, unlike in the covariate-shift setting, it is often the case in the spatial setting that test points are fixed rather than arising i.i.d. from a distribution. We first state our result and then discuss further intuition.
Proposition 4.3 (Counterexample to the consistency of 1NN).
For a particular spatial location $S$, let the associated test point be $(S,X,Y)$, with covariate $X$ as at the start of Section 2 and response $Y$ as in Assumption 2.2. Assume there exists a constant $c>0$ such that for all $S\in\mathcal{S}$ and associated $(S,X,Y)$, $\mathrm{Var}[\ell(Y,h^{\chi}(S))|S]\geq c$. Next, consider an infinite sequence of validation sets as in Def. 3.2. Suppose there exists an $S^{\prime}\in\mathcal{S}$ such that for any $r>0$ and $N^{\textup{{val}}}>0$, |$|\{1\leq j\leq N^{\textup{{val}}}:d_{\mathcal{S}}(S^{\textup{{val}}}_{j},S^{\prime})=r\}|\leq 1$. Choose any $Q^{\textup{{test}}}$ such that $Q^{\textup{{test}}}(\{S^{\prime}\})>0$. Then there exists a $\delta\in(3/4,1)$ and a $C^{(\delta)}>0$ such that, for each $N^{\textup{{val}}}$, with probability at least $1-\delta$, $|R^{\chi}_{Q^{\textup{{test}}}}(h)\!-\!(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|\geq C^{(\delta)}.$ Here $(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}$ denotes the 1NN estimator associated to the first $N^{\textup{{val}}}$ data points and $\delta$. $C^{(\delta)}$ and $\delta$ do not depend on $N^{\textup{{val}}}$ or other properties of the sequence of validation data.
See Section B.3.2 for a proof of the result. The technical condition |$|\{1\leq n\leq N^{\textup{{val}}}:d_{\mathcal{S}}(S^{\textup{{val}}}_{j},S^{\prime})=r\}|\leq 1$ ensures that the 1NN set for each point contains exactly $1$ point. That is, there are no ties. This condition would be satisfied for the infinite sequence of validation sets with probability one if, for instance, the validation points were chosen i.i.d. from a uniform measure on a compact set; cf. Prop. 3.3.
Intuition for Inconsistency of $1$-Nearest Neighbor. The variance of the 1NN estimator will, in non-trivial cases, remain non-negligible even as the amount of validation data increases without bound. Consider the simple case where $Q^{\textup{{test}}}=\delta_{S}$ for some $S\in\mathcal{S}$. Using $k=1$ would lead us to estimate the error using a single validation point, which is not consistent due to the noise in the observed response at the validation point. Where the problem with the holdout estimator was bias, the problem with 1NN is variance.
5 A Consistent Estimator
We next provide a novel bound on the test-risk estimation error of $k$-nearest neighbors. We propose using a $k$-nearest neighbors estimator with $k$ chosen by optimizing our bound. We show that our proposed estimator is consistent for test risk under infill asymptotics. Finally, we provide promising results for model selection but also discuss open challenges.
5.1 Our Bound and Estimator
Analogous to the intuition after Prop. 4.3, we expect inconsistency of $k$-nearest neighbors for any fixed choice of $k$. Instead, we propose to trade off the larger variance of small $k$ and larger bias of large $k$ by optimizing a bound depending on the validation set. To that end, we first derive a bound on the test-risk estimation error as a function of $k$.
[…]
$\zeta^{k}(\Psi_{1};\Psi_{2})=\sup_{S_{2}\in\Psi_{2}}\,\inf_{\begin{subarray}{c}A\subset\Psi_{1},\\ |A|=k\end{subarray}}\,\,\sup_{S_{1}\in A}d_{\mathcal{S}}(\Psi_{1},\Psi_{2}).$ (11)
Theorem 5.1 (Bound on Estimation Error in Terms of Fill Distance).
Consider a validation set $D^{\textup{{val}}}$ of size $N^{\textup{{val}}}$ and a test set $D^{\textup{{test}}}$ of size $M^{\textup{{test}}}$. Take the $k$-nearest neighbors test-risk estimator from Def. 4.2. Choose $k$ such that $1\leq k\leq N^{\textup{{val}}}$, $\delta\in(0,1)$, and $\rho_{k}\coloneqq\zeta^{k}(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},S^{\textup{{test}}}_{1:M^{\textup{{test}}}})$. Take Assumptions 2.2, 2.1, 2.4 and 2.5. Then, with probability at least $1-\delta$,
[…]
$\leq L\rho_{k}\!+\!\Delta\sqrt{\!\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\sqrt{\tfrac{1}{2}\log\tfrac{2}{\delta}},$ (13)
where $B(S,r)$ denotes the ball of radius $r$ centered at $S$. See Assumptions 2.1, 2.5 and 1 for $\Delta$,$L$, $Q^{\textup{{test}}}$ respectively.
We prove Thm. 5.1 in Section B.4. Roughly the first term on the righthand side in Eqns. 12 and 13 is a bias term, and the second term is a variance term. Specifically, the first term is a worst-case upper bound on the bias of our estimator; it will be large if the average loss varies quickly in space or if validation data is not available near test data. Larger $k$ may increase the first term. The second term comes from applying a tail bound; if most of the weight is put on a few validation points, the resulting estimator has high variance and this term is large. Sufficiently large $k$ will decrease the second term. If we can find a $k$ such that both (a) the distance from each test point to its $k$ nearest neighbors is small and (b) no validation point has too much impact on our estimator, $\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)$ provides a good estimate for $R^{\chi}_{Q^{\textup{{test}}}}$.
Thm. 5.1 is closely related to Portier et al. (2023, Prop. 4). While there are technical differences in the proof and algorithm (and the risk that is bounded), the substantive distinction is that we state our bound directly in terms of the fill distance, instead of upper bounding this distance again as done in Portier et al. (2023). We can therefore avoid making assumptions about the distributions of the sites; we instead highlight the fill distance of the validation set as an essential quantity in controlling the accuracy of nearest neighbor risk estimation.
Selection Procedure with Unknown Lipschitz Constant. If the Lipschitz constant of the average loss, $L$, can be upper bounded, for example by knowledge about how quickly varying the spatial processes involved in the analysis are, then $k$ can be selected by minimizing the upper bound in Eqn. 12. Since the bound is conditional on the validation and test sites, the bound still holds with the same probability for $k$ selected by this minimization. However, it will generally be the case that the Lipschitz constant is unknown. We therefore suggest choosing the number of neighbors by minimizing the upper bound from Thm. 5.1 with $1$ in place of the Lipschitz constant:
$k^{\star}_{T}\in\arg\min\nolimits_{k\in T}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2},$ (14)
with $\beta_{\delta}=\smash{\Delta\sqrt{\frac{1}{2}\log\frac{2}{\delta}}}$. For computational efficiency, we focus on choosing $k$ as a power of 2: $T=T_{2}:=\{2^{i}\}_{i=1}^{\smash{\lfloor\log_{2}N^{\textup{{val}}}\rfloor}}$. We call the resulting estimator spatial nearest neighbors (SNN).
5.2 Our Method is Consistent
We show that SNN is consistent under infill asymptotics.
Corollary 5.2 (Our Method is Consistent under Infill Asymptotics).
Suppose that $\mathcal{S}=[0,1]^{d}$. Take Assumptions 2.2, 2.1, 2.4 and 2.5. Let $k^{\star}_{T_{2}}\in\arg\min_{k\in T_{2}}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}$ with $\beta_{\delta}=\Delta\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$. Then $\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)$ is consistent under infill asymptotics.
See Section B.5 for a proof of Cor. 5.2. Cor. 5.2 tells us that minimizing an upper bound on our error in estimation leads to an estimator that is consistent regardless of the test data, as long as the validation data are dense on the unit cube.
5.3 Model Selection
A practitioner will commonly select a predictive method within some collection by choosing the method with lowest estimated test risk. We consider two cases.
(1) Fixed predictive method. So far we have focused on the setting where predictive methods are fixed in advance. As we accrue validation data, the consistency of SNN ensures it will eventually choose the predictive method with lowest test risk. Without consistency, holdout and 1NN cannot be trusted to choose the best predictive method.
(2) Proportional training and validation data. If we consider the special case where the predictive method is fit using training data (vs., say, a physical model), it is common for a practitioner to have a single set of available data that they then partition into training data and validation data; for instance, a fixed percentage of the total data may go to training. In this case, the predictive method changes as the validation set grows. Nonetheless, we can still be sure to eventually choose the predictive method with the lowest risk if the test-risk estimate has a faster rate of convergence in the number of validation data points than the convergence rate of the predictive method in the number of training points.
We are not able to formally characterize how SNN performs in model selection. But we provide rigorous results on rates of convergence of SNN that give suggestive guidance. In particular, (a) we first consider the case of a grid prediction task, where test sites are arranged on a grid. In this case, we are able to give a rate of convergence for both $1$NN and SNN (Cors. B.12 and B.13). If the validation data are i.i.d. our rate is faster than the minimax optimal convergence rate for predicting a Lipschitz function in the presence of additive, homoskedastic noise. Our result shows the same rate of convergence as Portier et al. (2023, Prop. 4), who considered test data that was i.i.d. instead of on a grid. Our result is suggestive that both nearest neighbor methods may perform well at model selection for validating maps.
(b) Second, we provide finite-sample bounds and asymptotic characterizations for general (non-grid) $Q^{\textup{{test}}}$ (Cor. B.10). In this case, when validation data is i.i.d., up to logarithmic factors, we show that SNN converges at the optimal rate of convergence for Lipschitz functions (Cor. B.14). Since in this case, the SNN convergence rate is not strictly faster than the training convergence rate, it might be difficult to select between statistical methods that converge at the optimal regression rate. But we would still expect to be able to select between (i) a statistical method that converges at the optimal rate and (ii) one that does not (for example a misspecified parametric model). Since the holdout and 1NN may not even be consistent, we could not rely on them to select the better model even in this latter, easier case.
6 Experiments
Figure 1: The validation sites (blue, clustered) for the first seed of the synthetic grid task. Panels from left to right and top to bottom represent $N^{\textup{{val}}}$ in the sequence $(250,500,1000,2000,4000,8000)$. Test sites (orange, gridded) are constant across panels.
We next provide experiments showing that the holdout and 1NN struggle in certain test-risk prediction tasks while SNN remains competitive across tasks. To be able to access ground truth, we mainly focus on simulations. But we also show the results of a real-data air temperature modelling task. In all experiments, we run SNN with the probability threshold $\delta=0.1$ and the loss upper bound $\Delta=1$.
6.1 Risk Estimation on Synthetic Data
We set up two synthetic experiments: a grid prediction task and a point prediction task. Based on our analyses above, we expect the holdout to struggle with the former and 1NN to struggle with the latter; our experiments confirm this intuition. See Section C.1.1 for full experiment details.
Figure 2: Error for test-risk estimation in the grid task across methods (holdout in blue, 1NN in orange, our SNN in green); lower values correspond to better performance. The vertical axis shows the absolute difference between the estimated test risk and empirical test risk. Each box plot shows the median, inter-quartile range, and outliers based on 100 synthetic datasets. The horizontal axis tracks increasing validation set sizes.
Validation and Test Data. In both experiments, we vary $N^{\textup{{val}}}\in\{250,500,1000,2000,4000,8000\}$. We use a truncated squared loss: $\ell(a,b)=\max((a-b)^{2},1.0)$. For the grid task, the test sites comprise a $50\times 50$ grid of equally spaced points in $[-0.5,0.5]^{2}$ (orange points in Fig. 1). We generate the validation sites via a sequential process that leads to clustering (blue points in Fig. 1). For the point task, there is a single test site at $(0,0)$. Validation sites are i.i.d. uniform in $[-0.5,0.5]^{2}$. For both tasks, we generate covariates and responses conditional on the sites:
$\!Y_{i}^{\smash{j}}$ $=f(\chi^{1}(S^{\smash{j}}_{i}),\chi^{2}(S^{\smash{j}}_{i}))+\eta(S^{j}_{i})+\epsilon_{i}^{\smash{j}},\,\epsilon^{j}_{i}\sim N(0,\sigma^{2})$
$\eta$ $\sim\mathcal{GP}(0,k_{1}),\,\chi^{1},\chi^{2}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}\mathcal{GP}(0,k_{2}),\,f\sim\mathcal{GP}(0,k_{3}),$
where $\mathcal{GP}(\mu,k)$ denotes a Gaussian process (GP) with mean function $\mu$ and kernel $k$. We describe our kernel and parameter choices in Section C.1.1. We plot examples of the generated data in Figs. 6 and 8.
Spatial Predictive Method. To arrive at our spatial predictive method, we generate training data according to the same distribution as the validation data. Since real-world analyses are often missing potentially relevant covariates, we retain only the first covariate (and not the second) as a realistic form of misspecification. We fit a GP regression, with zero prior mean and the same kernel used in data generation; we predict using the posterior mean.
Figure 3: Error for test-risk estimation in the point task across methods. Axes and box plot details are as in Fig. 2.
Ground Truth. We treat the empirical test risk $\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)$ as ground truth, with
$\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h):=(1/M^{\textup{{test}}})\sum\nolimits_{m=1}^{M^{\textup{{test}}}}\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))$ (15)
an unbiased Monte Carlo estimate of the test risk. To compute the empirical test risk, we take $2500$ samples of the covariates and responses at the test sites. [^7] For the point task, we generate 2500 $Y^{\textup{{test}}}_{i}$ values at the origin. By Hoeffding’s inequality, the empirical test risk is very close to the true test risk; namely, with probability $0.95$, $|\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)-R^{\chi}_{Q^{\textup{{test}}}}(h)|<0.028$. We expect even closer agreement in practice. Figs. 2 and 3 show the absolute difference between each estimator and the empirical test risk across 100 seeds.
Results. We expect our SNN estimator to be consistent in all tasks. In the grid task, we expect the variance of 1NN to be low since there are many, dense test points. And we expect the bias of the holdout to be high since the test and validation points have noticeably different spatial arrangement. Our results in the grid task (Fig. 2) agree with our intuitions; the errors of 1NN and our SNN decrease much more rapidly across the values of $N^{\textup{{val}}}$ than the holdout.
In the point task, we expect the high variance of 1NN to be an issue with substantial probability given that there is only one test point. Fig. 3 agrees with our intuition; the error of our SNN estimator decreases much more rapidly across the values of $N^{\textup{{val}}}$ than 1NN. In this case, we find that the holdout errors decrease rapidly as well. We also plot the (signed) relative difference of each estimator to the empirical test risk in Figs. 7 and 9.
In Tables 2 and 3, we show the number $k$ of nearest neighbors selected by SNN for each of the two tasks. In the grid task, the optimal $k$ was at most 2 in all cases we considered. In the point task, we found the optimal $k$ noticeably increased with $N^{\textup{{val}}}$, though it always remained over an order of magnitude smaller than $N^{\textup{{val}}}$.
Figure 4: Training (blue), validation (orange), and test (green) data for a single seed of the model selection experiment. The dashed red line depicts predictive method $h_{0}$, and dashed black shows $h_{1}$.
6.2 Model Selection on Synthetic Data
We next see that SNN and 1NN are able to select the model with lower test risk in a model selection task, but the holdout systematically picks the wrong model. We repeat the model selection problem 100 times. In each repetition, we have $N^{\text{train}}=100$ and a max $N^{\textup{{val}}}=75$. In our analysis, we will consider validation subsets of size $N^{\textup{{val}}}\in\{5\ell\}_{\ell=1}^{15}$. We generate independent test, validation, and training data as follows; see Fig. 4.
$U_{i}^{\smash{j}}$ $\sim\mathcal{U}([0,1]),\quad S_{i}^{\smash{j}}=\sqrt{U_{i}^{\smash{j}}},\quad j\in\{\text{train},\text{val}\}$
$S_{m}^{\smash{\text{test}}}$ $=m/20,0\leq m\leq 20,\quad\epsilon_{\smash{i}}^{\smash{j}}\sim\mathcal{U}([0,0.1]),$
$Y_{\smash{i}}^{j}$ $=|S_{\smash{i}}^{j}-0.5|+\epsilon_{\smash{i}}^{j}\quad j\in\{\text{train},\text{val},\text{test}\},$ (16)
We compare two predictive methods: $h_{0}(S)=0.25$ and $h_{1}(S)=\beta_{1}^{\top}S+\beta_{0}$, with $(\beta_{1},\beta_{0})$ fit by minimizing the mean absolute residual on the training data. Fig. 4 shows the data and predictions of both models (as functions of space). We use the loss function $\ell(a,b)=|a-b|$, which is bounded for this problem because both the hypotheses and the response variable are bounded on $[0,1]$.
Across all seeds, $h_{0}$ has the lower empirical test risk; $h_{1}$ makes large errors on the test points near 00 because most of the training data is in $[0.5,1]$. Since most of the validation data also clusters near 1, we expect the holdout to struggle due to bias. Our arguments in Section 5.3 lead us to expect both SNN and 1NN should perform well on this task when given sufficient validation data.
We say an estimator of the risk, $\hat{R}$, selects $h_{0}$ if $\hat{R}(h_{0})<\hat{R}(h_{1})$. We plot the percentage of times each method (correctly) selects $h_{0}$ as a function of the number of validation points in Fig. 5. When the validation set is small, all estimators select the model with lowest test risk ($h_{0}$) less than half the time. For the nearest neighbor methods, we expect that when there are few or no spatial locations less than $1$, weighting cannot fix the estimate. However, when the number of validation points is large, the nearest neighbor methods consistently (correctly) select $h_{0}$. By contrast, the holdout consistently (incorrectly) selects $h_{1}$, even though $h_{1}$ has higher test risk. See Section C.2 for full experiment details.
Figure 5: The percentage of times each estimator (correctly) selects the model with lower empirical test risk ($h_{0}$), out of 100 seeds as a function of $N^{\textup{{val}}}$. Estimators include the holdout (blue), 1NN (orange), and our SNN (green).
6.3 Air Temperature Prediction
Although we do not have access to ground truth in real data, we next show a real-data experiment where the holdout, 1NN, and our SNN give very different estimates of test risk and can differ in model selection.
We consider predicting monthly average air surface temperature based on remote sensing observations of monthly average land surface temperature at day and night from MODIS-Aqua (Wan et al., 2021). This task is inspired by Hooker et al. (2018), who used two weighted regression methods to construct a global map of air temperature.
Our test task is prediction at the 5 largest urban areas in the United States (New York City, Los Angeles, Chicago, Miami, and Houston). Loss is absolute error in degrees Celsius. For predictive methods, we consider geographically weighted linear regression (GWR) and a baseline GP spatial interpolation model fit only to weather station data.
Table 1 shows a large discrepancy in test-risk estimates across estimators. We see that 1NN chooses a different predictive method (spatial GP) than holdout or our SNN does. Given the results in the rest of this paper, we would favor decisions made with SNN rather than the other two estimators. See Section C.3 for full details and also a grid-prediction task where all estimators are in agreement.
Table 1: Test-risk estimates for the 5-metros task. Rows correspond to predictive methods and columns to estimators. We report two standard error intervals for holdout. For each estimator, we bold the predictive method with lower estimated test risk.
Holdout 1NN SNN
GWR $\mathbf{0.84\pm 0.03}$ $0.60$ $\mathbf{0.67}$
Spatial GP $0.96\pm 0.04$ $\mathbf{0.40}$ $0.75$
7 Discussion
We proposed an adaptive nearest neighbors method for predictive validation. We showed that past validation methods can be inconsistent under infill asymptotics but our method is consistent. A number of open questions remain for spatial validation, such as providing checks when an infill assumption is inappropriate or checks on extrapolation quality.
Acknowledgements
This work was supported in part by an NSF CAREER Award and the Office of Naval Research under grant N00014-20-1-2023 (MURI ML-SCOPE).
Appendix A Extended Related Work
[…]
A.1 Validation versus Cross-Validation
While we expect our work to have implications for cross validation (CV), we focus on validation here since (1) we do not assume training data exists or is easy to change and (2) CV presents additional subtle challenges. The holdout is broadly applicable to any predictive method (whether data-driven, physical, or a combination thereof). However, in the cases when data is scarce and a data-driven predictive method is used, cross-validation is commonly believed to make more efficient use of the data. CV is widely used in spatial analyses; among many examples are Wang et al. (2022); Valavi et al. (2021); Kianian et al. (2021). However, the interpretation of the error estimates given by cross-validation is subtle even in the classical i.i.d. setting (Bates et al., 2023). We focus on the validation setting in this work due to its broad applicability and clear interpretation, though extensions to cross-validation, as well as a theoretical understanding of how the resulting error estimate relate to predictive performance is a promising direction for future work.
A.2 Covariate Shift
In covariate shift, it is generally assumed that $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P^{\text{val}}$, $S^{\textup{{test}}}_{m}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}Q$ and $U^{\ell}_{i}|S^{\ell}_{i}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}\mu$ for $\ell\in\{\text{val},\text{test}\}$. Moreover, it is typically assumed the density ratio $\frac{dQ}{dP}$ exists and is bounded, although there is recent work relaxing this assumption in the context of non-parametric regression (Kpotufe & Martinet, 2021; Pathak et al., 2022). Mean estimation seeks to estimate $\mathbb{E}[U^{\text{test}}_{1}]$. Taking $(U^{\text{test}}_{m}|S^{\textup{{test}}}_{m})=\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))$, this is the same task we consider, but with the (stronger) assumption that covariates are independent and identically distributed with a bounded density ratio between the test and validation distributions. This assumptions is not appropriate for validation with many spatial datasets, with particularly simple examples being when the task of interest requires prediction at a single spatial location, or on a regular grid. Our assumptions are essentially a conditional formulation of mean estimation under covariate shift. Many methods proposed for addressing mean estimation under covariate shift, including all the approaches we describe below, are based on re-weighting validation points, then applying the holdout approach described earlier with these weights.
The kernel mean matching algorithm provides a solution under standard covariate shift assumptions with a bounded density ratio and assuming conditional expectation $e^{\chi}_{h}$ lives in a reproducing kernel Hilbert space (RKHS) and has small norm (Gretton et al., 2009). Yu & Szepesvári (2012) provide a finite sample bound for this method
Appendix B Proofs of Claims
We now present results and proofs not included in the main text. We essentially follow the order of results in the main text. Section Section B.1 gives sufficient conditions for Assumption 2.5 to hold for a homoskedastic, additive noise model and squared loss for responses taking values in $[0,1]$, mentioned in Section 2.2. Section B.2 focuses on the proof of Prop. 3.3, first recalling several properties of covering numbers that will be used in the result. In Section B.3 we restate and prove our Prop. 4.1 and Prop. 4.3, which show limitations of existing validation methods in the spatial setting we study. In Section B.4 we prove a general result upper bounding the error in estimating the risk using $k$-nearest neighbors, as well as an upper bound on the error for $k^{\star}_{T_{2}}$. Section B.5 establishes the consistency of spatial nearest neighbors. Section C.2 discusses issues related to model selection and proves rates of convergence for spatial nearest neighbors.
B.1 Lipschitz Constant for Lipschitz Response and Predictive Method
While assuming the average loss is Lipschitz continuous as a function of space is mathematically convenient, it is perhaps more natural to make assumptions about the spatial field we are trying to make predictions about, as well as the predictive method we are using to make prediction. The following proposition gives an example of how Lipschitz continuity of the processes involved can imply Assumption 2.5.
Proposition B.1.
Consider $\mathcal{Y}\subset[0,1]$ and squared loss. Then if $f^{\chi}(S):=f(S,\chi(S))$ is $L_{Y}$-Lipschitz and $h^{\chi}$ is $L_{h}$-Lipschitz. Let $(S,X,Y)$ be generated as in Assumption 2.2, with $Y=f(S,X)+\epsilon$. Then $e^{\chi}_{h}(S)\coloneqq\mathbb{E}[(Y-h^{\chi}(S))^{2}|S]$ is $2(L_{Y}+L_{h})$-Lipschitz.
Proof.
Let $S,S^{\prime}\in\mathcal{S}$ and $\epsilon,\epsilon^{\prime}$ be the associated noise random variables. Then,
$|e^{\chi}_{h}(S)-e^{\chi}_{h}(S^{\prime})|$ $=\mathbb{E}[(f^{\chi}(S)+\epsilon)-h^{\chi}(S))^{2}-(f^{\chi}(S^{\prime})+\epsilon^{\prime})-h^{\chi}(S^{\prime}))^{2}]$ (17)
$=f^{\chi}(S)^{2}-f^{\chi}(S)^{2}+h^{\chi}(S)^{2}-h^{\chi}(S^{\prime})^{2},$ (18)
where we have used that $\mathbb{E}[\epsilon]=\mathbb{E}[\epsilon^{\prime}]=0$, $\epsilon,\epsilon^{\prime}$ are independent from $h$ and $\mathbb{E}[\epsilon^{2}]=\mathbb{E}[(\epsilon^{\prime})^{2}]$. Then,
$f^{\chi}(S)^{2}-f^{\chi}(S^{\prime})^{2}+h^{\chi}(S)^{2}-h^{\chi}(S^{\prime})^{2}$ $=(f^{\chi}(S)+f^{\chi}(S^{\prime}))(f^{\chi}(S)-f^{\chi}(S^{\prime}))+(h^{\chi}(S)+h^{\chi}(S^{\prime}))(h^{\chi}(S)-h^{\chi}(S^{\prime}))$
$\leq 2|f^{\chi}(S)-f^{\chi}(S^{\prime})|+2|h^{\chi}(S)-h^{\chi}(S^{\prime})|$ (19)
$\leq 2(L_{Y}+L_{h})d_{\mathcal{S}}(S,S^{\prime}).$ (20)
The first inequality uses that $\mathcal{Y}\subset[0,1]$ and the second the Lipschitz assumptions on $f$ and $h$. ∎
Therefore, at least in the case of squared loss with bounded response and predictive method values and a homoskedastic additive noise model, smoothness of the average response surface (as a function of space) together with smoothness of the predictive method imply Assumption 2.5.
B.2 Proof that Independent and Identically Distributed Data Implies Infill Asymptotic with High Probability
The purpose of this section is to prove Prop. 3.3. We begin by recalling this proposition:
See 3.3
Our proof is similar to earlier proofs in Reznikov & Saff (2015, Section 5.2) and essentially follows the stackexchange response (2021) (https://mathoverflow.net/users/36721/iosif pinelis) keeping track of numerical constants. Essentially, the idea is that were the fill distance to be large, there must be a ball of large radius that doesn’t contain any points in the sample of location. But because the probability of a point in the sample falling in any ball is bounded below in terms of the volume of the ball, this must be improbable as the sample size grows.
B.2.1 Preliminary Definitions and Lemmas Related to Nets and Covering Number
In order to formalize the proof of Prop. 3.3 sketched above idea, we recall the definition of a net and covering number, as well as some standard properties of covering numbers.
Definition B.2 (Net, Covering Number).
Let $A\subset\mathbb{R}^{d}$ compact. Any finite set $C\subset\mathbb{R}^{d}$ such that
$A\subset\bigcup_{S\in C}B(S,\epsilon)$ (21)
where $B(S,\epsilon)$ denotes the $d$-dimensional closed Euclidean ball of radius $\epsilon\geq 0$ centered at $S$ is referred to as an $\epsilon$-net of $A$. The $\epsilon$-covering number of $A$, denoted by $\mathfrak{N}(\epsilon,A)$ is the minimum cardinality of an $\epsilon$-net of $A$.
Because $A$ is compact for any $\epsilon$ the $\epsilon$-covering number of $A$ is finite, and will generally increase as $\epsilon\to 0$.
We now recall how covering number behaves under affine transformations of the underlying space. For set $A,B$, we define $A+B=\{a+b:a\in A,b\in B\}$.
Proposition B.3.
For any compact $A\subset\mathbb{R}^{d}$ and $c\in\mathbb{R}^{d}$, $\mathfrak{N}(\epsilon,A)=\mathfrak{N}(\epsilon,A+\{c\})$. For any $\alpha>0$, $\mathfrak{N}(\epsilon,A)=\mathfrak{N}(\alpha\epsilon,\alpha A)$.
Proof.
The first claim is shown by noting that if $C$ is an $\epsilon$-net of $A$ then $C+\{c\}$ is an $\epsilon$-net of $A+\{c\}$ so that $\mathfrak{N}(\epsilon,A+\{c\})\leq\mathfrak{N}(\epsilon,A)$. Applying a symmetric argument implies the reverse inequality.
For the second claim, let $C$ be an $\epsilon$-net for $A$. Then,
$\alpha A\subset\alpha\left(\bigcup_{S\in C}B(S,\epsilon)\right)=\bigcup_{S\in C}B(\alpha S,\alpha\epsilon)=\bigcup_{S^{\prime}\in\alpha C}B(S^{\prime},\alpha\epsilon),$ (22)
and so $\alpha C$ is an $\alpha\epsilon$-net of $\alpha A$. It follows that $\mathfrak{N}(\alpha\epsilon,\alpha A)\leq\mathfrak{N}(\epsilon,A)$. The opposite inequality is obtained via the same argument applied with $\alpha^{\prime}=\frac{1}{\alpha},$ $A^{\prime}=\alpha A$ and $\epsilon^{\prime}=\alpha\epsilon$. ∎
In the proof of Prop. 3.3 we apply a union bound over all elements in a net for the unit cube. We therefore need a result telling us that this net does not contain too many elements.
Lemma B.4 (Covering number of Unit Cube).
Let $\epsilon\in(0,1]$. Then $\mathfrak{N}(\epsilon,[-1,1]^{d})\leq\frac{1}{B_{d}}\left(\frac{6}{\epsilon}\right)^{d}$
where $B_{d}=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2}+1)}$ is the volume of the $d$-dimensional unit sphere.
Proof.
By the upper bound in Wainwright (2019, Lemma 5.7),
$\mathfrak{N}(\epsilon,[-1,1]^{d})\leq\frac{1}{B_{d}}\text{vol}\left(\left[-\frac{2}{\epsilon},\frac{2}{\epsilon}\right]^{d}+B(0,1)\right).$ (23)
For any $S=S_{1}+S_{2}\in[-\frac{2}{\epsilon},\frac{2}{\epsilon}]^{d}+B(0,1)$ with $S_{1}\in[-\frac{2}{\epsilon},\frac{2}{\epsilon}]^{d}$ and $S_{2}\in B(0,1)$, by the triangle inequality for infinity norm,
$\|S\|_{\infty}\leq\frac{2}{\epsilon}+1\leq\frac{3}{\epsilon},$ (24)
where we first used that points in the unit ball have infinity norm not more than $1$ and then used that $\epsilon\leq 1$. Therefore,
$\text{vol}\left(\left[-\frac{2}{\epsilon},\frac{2}{\epsilon}\right]^{d}+B(0,1)\right)\leq\text{vol}\left(\left[-\frac{3}{\epsilon},\frac{3}{\epsilon}\right]^{d}\right)=\left(\frac{6}{\epsilon}\right)^{d}.$ (25)
B.2.2 Main Proof
We recall and prove Prop. 3.3.
The proof essentially follows (2021) (https://mathoverflow.net/users/36721/iosif pinelis), making constants explicit. The idea of the proof is that the distance from any point in $[0,1]^{d}$ can’t be larger than the distance from this point to the nearest point in some net, plus the distance from that point to the nearest point in the validation sample. Therefore, it suffices to show that at least one point in the sample falls in a ball of some radius around each point in a net. This is done with a union bound over all elements in a minimal net, using that the sampling density of validation points is bounded below and so the probability of any particular validation point falling in a ball centered at any point in the net can’t be too small.
We again recall Prop. 3.3 for convenience when reading the proof. See 3.3
Proof of Prop. 3.3.
Fix $\tau\in(0,1)$ and let $C$ be a minimal cardinality $\tau/2$-net for $[0,1]^{d}$. If $B(S,\tau/2)$ contains a validation point for all $S\in C$, then for any $S^{\prime}\in[0,1]^{d}$, by the triangle inequality,
$\min_{1\leq n\leq N^{\textup{{val}}}}d_{\mathcal{S}}(S^{\prime},S^{\textup{{val}}}_{n})\leq\min_{1\leq n\leq N^{\textup{{val}}}}\left(\min_{S^{\prime}\in C}d_{\mathcal{S}}(S^{\prime},S)+d_{\mathcal{S}}(S^{\textup{{val}}}_{n},S)\right)\leq\tau.$ (26)
Therefore the probability that the fill distance is large $(>\tau)$ is less than the probability that there exists an element of the net such that no validation point is close to it (within radius $\tau/2$):
$\Pr(\zeta(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},[0,1]^{d})>\tau)\leq\Pr(\exists S\in C:S^{\textup{{val}}}_{n}\not\in B(S,\tau/2)\,\forall 1\leq n\leq N^{\textup{{val}}}).$ (27)
The probability of any particular validation point falling in a ball centered at any point of radius contained in the $\tau/2$ can’t be too small since $P$ has density that is bounded below: For any $S\in[0,1]^{d}$
[…]
where $\lambda$ denotes Lebesgue measure and in the final inequality we have used that since $S\in[0,1]^{d}$, at least one quadrant of $B(S,\tau/2)$ is contained in $[0,1]^{d}$.
Returning to Eqn. 27 and taking a union over all elements in the net:
$\Pr(\zeta(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},[0,1]^{d})>\tau)$ $\leq\mathfrak{N}(\tau/2,[0,1]^{d})\max_{a\in A}\Pr(S^{\textup{{val}}}_{n}\not\in B(a,\tau/2)\,\forall 1\leq n\leq N^{\textup{{val}}})$ (29)
$\leq\mathfrak{N}(\tau/2,[0,1]^{d})\max_{a\in A}(1-P(B(a,\tau/2)))^{N^{\textup{{val}}}}$ (30)
$\leq\mathfrak{N}(\tau/2,[0,1]^{d})\left(1-\frac{c}{2^{d}}B_{d}\left(\frac{\tau}{2}\right)^{d}\right)^{N^{\textup{{val}}}},$ (31)
where $B_{d}=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2}+1)}$ is the volume of the $d$-dimensional unit sphere. Eqn. 30 uses Eqn. 28. We now upper bound the terms in Eqn. 31.
Applying, Props. B.3 and B.4
$\mathfrak{N}(\tau/2,[0,1]^{d})\leq\frac{1}{B_{d}}\left(\frac{6}{\tau}\right)^{d}.$ (32)
Using the inequality $(1-x)\leq e^{-x}$,
$\left(1-\frac{c}{2^{d}}B_{d}\left(\frac{\tau}{2}\right)^{d}\right)^{N^{\textup{{val}}}}\leq\exp\left(-\frac{c}{2^{d}}{N^{\textup{{val}}}}B_{d}\left(\frac{\tau}{2}\right)^{d}\right).$ (33)
Combining Eqns. 31, 32 and 33,
$\Pr(\zeta(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},[0,1]^{d})>\tau)\leq\frac{1}{B_{d}}\left(\frac{6}{\tau}\right)^{d}\exp\left(-\frac{c}{2^{d}}{N^{\textup{{val}}}}B_{d}\left(\frac{\tau}{2}\right)^{d}\right)$ (34)
Now choose $\tau^{d}=\frac{4^{d}}{cN^{\textup{{val}}}B_{d}}\log\frac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}$. For all $N^{\textup{{val}}}$ larger than some $n_{0}$, $\tau<1$ because $\lim_{N^{\textup{{val}}}\to\infty}\frac{4^{d}}{cN^{\textup{{val}}}B_{d}}\log\frac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}=0$. For this choice of $\tau$
$\frac{1}{B_{d}}\left(\frac{6}{\tau}\right)^{d}\exp\left(-\frac{c}{2^{d}}{N^{\textup{{val}}}}B_{d}\left(\frac{\tau}{2}\right)^{d}\right)=\frac{1}{\tau^{d}N^{\textup{{val}}}}\delta$ $=\frac{cB_{d}\delta}{4^{d}\log\frac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}}\leq\delta\frac{c}{2^{d}\log\frac{3^{d}N^{\textup{{val}}}}{\delta}},$ (35)
where in the final inequality we use that $B_{d}\leq 2^{d}$ since the unit ball is contained in the unit cube for $\tau<1$. For all $N^{\textup{{val}}}\geq\frac{\delta e^{c/2^{d}}}{3^{d}}$ the right-hand side is less than $\delta$; but $\delta<1,c\leq 1$ and so this holds for all $N^{\textup{{val}}}\geq 1$. ∎
B.3 Proof of Inconsistency of Existing Methods
In this section, we restate and prove our results on limitations of existing methods for validation. These results were stated in Section 4 in the main text.
B.3.1 Inconsistency of the Holdout
We begin by focusing on the holdout. As sketched in the main text, the holdout does not depend on the particular test set, and therefore cannot approximate the test risk well for point prediction tasks unless the average loss at both points is the same.
See 4.1
Proof.
Because $e^{\chi}_{h}(S)$ is not a constant $\mathcal{S}$ contains at least two elements $S,S^{\prime}$ such that $e^{\chi}_{h}(S)\neq e^{\chi}_{h}(S^{\prime})$. Let $Q=\delta_{S}$ and $Q^{\prime}=\delta_{S^{\prime}}$. There there exists some $\gamma>0$ such that
$|R^{\chi}_{Q}(h)-R^{\chi}_{Q^{\prime}}(h)|=|e^{\chi}_{h}(S)-e^{\chi}_{h}(S^{\prime})|>2\gamma.$ (36)
If $\hat{R}^{\chi}_{\textup{{Hold}}}$ is not consistent for $R^{\chi}_{Q}(h)$, we are done. Otherwise, by the definition of consistency, for all $N\geq N_{0}$ with probability at least $1/2$,
$|(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}(h)-R^{\chi}_{Q}(h)|<\gamma.$ (37)
where we use $(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}$ to denote the estimator constructed using the first $N$ validation points, $D^{\textup{{val}}}_{N}$. By the reverse triangle inequality,
$|(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}(h)-R^{\chi}_{Q^{\prime}}(h)|\geq|R^{\chi}_{Q}(h)-R^{\chi}_{Q^{\prime}}(h)|-|(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}(h)-R^{\chi}_{Q^{\prime}}(h)|.$ (38)
Combining Eqn. 37 and Eqn. 38, for all $N\geq N_{0}$ with probability at least $1/2$
$|(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}(h)-R^{\chi}_{Q^{\prime}}(h)|>2\gamma-\gamma=\gamma,$ (39)
which implies the holdout is not consistent for $R^{\chi}_{Q^{\prime}}(h)$. ∎
B.3.2 Inconsistency of $1$-nearest neighbor Estimator
We now turn to $1$-nearest neighbor risk estimation and restate and prove Prop. 4.3: See 4.3
The idea of the proof is that if the loss has a non-vanishing variance, then the one nearest neighbor procedure results in an estimator with a non-zero variance for point prediction. Therefore, it cannot converge in probability to the actual risk, which is deterministic.
B.3.3 Preliminary Result
We begin by proving a result that says that if a bounded random variable has a second moment bounded below by $C$, then it cannot be close to zero most of the time. We will apply this inequality to the second moment of the difference between the $1$-nearest neighbor estimator and the test risk to in our proof of Prop. 4.3.
Proposition B.5.
Let $U$ be a random variable with $U\in[-A,A]$ almost surely and $\mathbb{E}[U^{2}]\geq C>0$. Then for any $\delta\in(1-\frac{C}{A^{2}},1)$ with probability $1-\delta$
$|U|\geq A\sqrt{1-\frac{1-\frac{C}{A^{2}}}{\delta}}.$ (40)
Proof.
Because $U\in[-A,A]$, $A^{2}-U^{2}$ is a non-negative random variable. Applying Markov’s inequality, for any $t\geq 0$,
$\mathrm{Pr}(A^{2}-U^{2}\geq t)\leq\frac{A^{2}-\mathbb{E}[U^{2}]}{t}\leq\frac{A^{2}-C}{t}.$ (41)
Take $t=\frac{A^{2}-C}{\delta}$ which is greater than 00 because $\delta\in(1-\frac{C}{A^{2}},1)$. Then Eqn. 41 becomes,
$\mathrm{Pr}\left(A^{2}-U^{2}\geq\frac{A^{2}-C}{\delta}\right)\leq\delta.$ (42)
Taking complements, with probability at least $1-\delta$
$A^{2}-U^{2}<\frac{A^{2}-C}{\delta}.$ (43)
Rearranging implies that with probability at least $1-\delta$
$U^{2}>A^{2}-\frac{A^{2}-C}{\delta}.$ (44)
For $\delta\in(1-\frac{C}{A^{2}},1)$ this bound is non-vacuous (strictly greater than zero). Taking square roots, which is monotone, for any such $\delta$ with probability at least $1-\delta$
$|U|>A\sqrt{1-\frac{1-\frac{C}{A^{2}}}{\delta}}.$ (45)
B.3.4 Restatement and Proof of Inconsistency of $1$-nearest neighbor
We return to our proof of Prop. 4.3. The idea will be to consider a point prediction task and then apply Prop. B.5 to show that with some fixed probability, the $1$-nearest neighbor estimator is a fixed distance away from the test risk, even as $N^{\textup{{val}}}$ increases.
See 4.3
Proof.
Because expectation minimizes the squared error to a random variable over all constant functions
$\mathbb{E}|R^{\chi}_{Q^{\textup{{test}}}}(h)-(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|^{2}\geq\mathbb{E}|(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)-\mathbb{E}[(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)]|^{2}.$ (46)
Because the $\epsilon^{\textup{val}}_{n}$ are independent the variance of $(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}$ is additive
$\mathbb{E}|(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)-\mathbb{E}(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|^{2}=\sum_{n=1}^{N^{\textup{{val}}}}(w^{\textup{{NN}},\!1}_{n})^{2}\mathbb{E}[\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n}))-e^{\chi}_{h}(S^{\textup{{val}}}_{n})]\geq V\sum_{n=1}^{N^{\textup{{val}}}}(w^{\textup{{NN}},\!1}_{n})^{2}.$ (47)
where $40<V<\Delta^{2}/$ 4 is the assumed lower bound on the variance of $\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n}))-e^{\chi}_{h}(S^{\textup{{val}}}_{n})$ and we have left implicit the dependence of the weights on $N^{\textup{{val}}}$. Also, |$|\{1\leq n\leq N^{\textup{{val}}}:d_{\mathcal{S}}(S^{\prime},S^{\textup{{val}}}_{n})=r|\leq 1$, implies $S^{\prime}$ has exactly one $1$-nearest neighbor in $S^{\textup{{val}}}_{1:N^{\textup{{val}}}}$, call the index of this neighbor $n(S^{\prime})$. Then
$\sum_{n=1}^{N^{\textup{{val}}}}(w^{\textup{{NN}},\!1}_{n})^{2}\geq(w^{\textup{{NN}},\!1}_{n(S^{\prime})})^{2}\geq Q^{\textup{{test}}}(\{S^{\prime}\})^{2}.$ (48)
Combining Eqn. 47 and Eqn. 48
$\mathbb{E}[|R^{\chi}_{Q^{\textup{{test}}}}(h)-(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|^{2}]\geq VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}.$ (49)
We now apply Prop. B.5 with $U=|R^{\chi}_{Q^{\textup{{test}}}}(h)-(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|$ to conclude that for $\delta\in(1-\frac{VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}}{\Delta^{2}},1)$ with probability at least $1-\delta$
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|\geq\Delta\sqrt{1-\frac{1-\frac{VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}}{\Delta^{2}}}{\delta}}>0.$ (50)
As neither $\delta$ nor the right hand side of Eqn. 50 depend on $N^{\textup{{val}}}$, $1$-nearest neighbor is not consistent under infill asymptotics. ∎
B.4 General Nearest Neighbor Bound and Selecting the Number of Neighbors
In this section, we prove two results giving bounds on the performance of nearest neigbhor risk estimation that will be the basis of later results. The first, already stated in the main text, is a general bound for any $k$.
See 5.1
The second result we will show is specific to SNN and relates the error incurred by using $k^{\star}_{T_{2}}$ to the error of the minimizer of the bound from Thm. 5.1:
nnminimizationpowtwo
B.4.1 Preliminary Result: Hoeffding’s Inequality
We begin by recalling Hoeffding’s inequality, which will be used to control tail probabilities of the sum of the weighted losses being far from its expectation.
Lemma B.6 (Hoeffding’s Inequality, (Hoeffding, 1963, Theorem 2)).
Let $(Z_{i})_{i=1}^{\ell}$ be independent random variables and $(a_{i})_{i=1}^{\ell}$ and $(b_{i})_{i=1}^{\ell}$ be sequences of real numbers such that $a_{i}\leq Z_{i}\leq b_{i}$ almost surely. Then for all $t>0$
$\mathrm{Pr}\left(\left\lvert\sum_{i=1}^{\ell}Z_{i}-\sum_{i=1}^{\ell}\mathbb{E}[Z_{i}]\right\rvert\geq t\right)\leq 2\exp\left(-\frac{2t^{2}}{\sum_{i=1}^{\ell}(b_{i}-a_{i})^{2}}\right).$ (51)
Equivalently, for any $\delta\in(0,1)$ with probability at least $1-\delta$
$\left\lvert\sum_{i=1}^{\ell}Z_{i}-\sum_{i=1}^{\ell}\mathbb{E}[Z_{i}]\right\rvert\leq\|b-a\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}.$ (52)
where $a,b\in\mathbb{R}^{\ell}$ have entries $a_{i}$ and $b_{i}$ respectively.
B.4.2 Proof of General Nearest Neighbor Risk Estimation Bound
We again recall and prove Thm. 5.1. The idea is to apply triangle inequality to split the error into a bias term and a sampling error term. The sampling error term is then controlled with Lemma B.6 since the loss is bounded. The bias term is controlled using Assumption 2.5. See 5.1
Proof.
By the triangle inequality,
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)|$ $\leq\underbrace{\big{|}R^{\chi}_{Q^{\textup{{test}}}}(h)-\sum_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}e^{\chi}_{h}(S^{\textup{{val}}}_{n})\big{|}}_{\tau_{1}}+\underbrace{\Big{|}\sum_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}(\ell(f(S^{\textup{{val}}}_{n},\chi(S^{\textup{{val}}}_{n}),\epsilon^{\textup{val}}_{n}),h^{\chi}(S^{\textup{{val}}}_{n}))-e^{\chi}_{h}(S^{\textup{{val}}}_{n})\Big{|}}_{\tau_{2}}.$ (53)
The first term, $\tau_{1}$ is a bias term, while the second term, $\tau_{2}$ is a sum of $N^{\textup{{val}}}$ independent variables with expectation zero. Using Assumption 2.1, we apply Hoeffding’s inequality (Lemma B.6) to bound $\tau_{2}$: for any $\delta\in(0,1)$ with probability at least $1-\delta$
$\tau_{2}\leq\Delta\|w^{\textup{{NN}},k}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}.$ (54)
By Hölder’s inequality and because the weights are non-negative and sum to one,
$\|w^{\textup{{NN}},k}\|_{2}\leq\sqrt{\|w^{\textup{{NN}},k}\|_{1}\|w^{\textup{{NN}},k}\|_{\infty}}=\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}}.$ (55)
Recalling the definition of $w^{\textup{{NN}},k}$ (Def. 4.2) and that each $A^{k}(s)$ contains at least $k$ points by construction,
$w^{\textup{{NN}},k}_{n}=\frac{1}{M^{\textup{{test}}}}\sum_{m=1}^{M^{\textup{{test}}}}\frac{1}{|A^{k}(S^{\textup{{test}}}_{m})|}\mathbf{1}\{S^{\textup{{val}}}_{n}\in A^{k}(S^{\textup{{test}}}_{m})\}\leq\frac{1}{k}\frac{1}{M^{\textup{{test}}}}\sum_{m=1}^{M^{\textup{{test}}}}\mathbf{1}\{S^{\textup{{val}}}_{n}\in A^{k}(S^{\textup{{test}}}_{m})\}.$ (56)
By the definition of $\rho_{k}$,
$S^{\textup{{val}}}_{n}\in A^{k}(S^{\textup{{test}}}_{m})\Rightarrow S^{\textup{{test}}}_{m}\in B(S^{\textup{{val}}}_{n},\rho_{k})$ (57)
and so
$\mathbf{1}\{S^{\textup{{val}}}_{n}\in A^{k}(S^{\textup{{test}}}_{m})\}\leq\mathbf{1}\{S^{\textup{{test}}}_{m}\in B(S^{\textup{{val}}}_{n},\rho_{k})\}.$ (58)
Substituting this in Eqn. 56 and using Eqn. 55
$\|w^{\textup{{NN}},k}\|_{2}\leq\sqrt{\frac{1}{k}\max_{1\leq n\leq N^{\textup{{val}}}}\frac{1}{M^{\textup{{test}}}}\sum_{m=1}^{M^{\textup{{test}}}}\mathbf{1}\{S^{\textup{{test}}}_{m}\in B(S^{\textup{{val}}}_{n},\rho_{k})\}}=\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}.$ (59)
It remains to bound the bias term, $\tau_{1}$. Define $\alpha_{nm}^{k}=\frac{1}{|A^{k}(S^{\textup{{test}}}_{m})|}\mathbf{1}\{S^{\textup{{val}}}_{n}\in A^{k}(S^{\textup{{test}}}_{m})\}$. Recalling the definition of $w^{\textup{{NN}},k}$ and rearranging the order of summation
$\tau_{1}=\left\lvert\frac{1}{M^{\textup{{test}}}}\sum_{m=1}^{M^{\textup{{test}}}}(e^{\chi}_{h}(S^{\textup{{test}}}_{m})-\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{nm}^{k}e^{\chi}_{h}(S^{\textup{{val}}}_{n}))\right\rvert\leq\max_{1\leq m\leq M^{\textup{{test}}}}\left\lvert e^{\chi}_{h}(S^{\textup{{test}}}_{m})-\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{nm}^{k}e^{\chi}_{h}(S^{\textup{{val}}}_{n})\right\rvert.$ (60)
Because for any $1\leq m\leq M^{\textup{{test}}}$, $\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{nm}^{k}=1$
$\left\lvert e^{\chi}_{h}(S^{\textup{{test}}}_{m})-\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{mn}^{k}e^{\chi}_{h}(S^{\textup{{val}}}_{n})\right\rvert=\left\lvert\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{mn}^{k}(e^{\chi}_{h}(S^{\textup{{test}}}_{m})-e^{\chi}_{h}(S^{\textup{{val}}}_{n}))\right\rvert\leq\max_{n:\alpha_{mn}^{k}>0}|e^{\chi}_{h}(S^{\textup{{test}}}_{m})-e^{\chi}_{h}(S^{\textup{{val}}}_{n})|.$ (61)
Applying Assumption 2.5 and taking the maximum over $m$ as well,
$\tau_{1}\leq\max_{n,m:\alpha_{nm}^{k}>0}Ld_{\mathcal{S}}(S^{\textup{{test}}}_{m},S^{\textup{{val}}}_{n}).$ (62)
The constraint $\alpha_{nm}^{k}>0$ implies that $d_{\mathcal{S}}(S^{\textup{{test}}}_{m},S^{\textup{{val}}}_{n})\leq\rho_{k}$ and so
$\tau_{1}\leq L\rho_{k}.$ (63)
The result follows from combining Eqns. 53, 54, 63 and 59. ∎
B.4.3 Proofs Related to Selecting the Number of Neighbors
We now restate and prove a bound that upper bounds the error of risk estimation with SNN (i.e. using $k^{\star}_{T_{2}}$ neighbors) to the minimum of the upper bound from Thm. 5.1 over all $k$. A key observation is that because Thm. 5.1 is conditional on the test locations, it can be minimized without the need to take a union bound over all $k$ in the set we minimize over. We first need a preliminary result, which holds for minimization over any subset of $\{1,\dots,N^{\textup{{val}}}\}$.
Proposition B.7 (Minimization of Upper Bound).
Let $T\subset\{1,\dots,N^{\textup{{val}}}\}$. Fix $\delta\in(0,1)$. Define $k^{\star}_{T}\in\arg\min_{k\in T}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}$ with $\beta_{\delta}=\Delta\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$. Define $C_{L}=\max(1,L)$. Under Assumptions 2.2, 2.1 and 2.5 with probability at least $1-\delta$
||| $R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T}}}(h)|\leq C_{L}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}\right)$ (64)
$\leq C_{L}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\right),$ (65)
where $B(S,r)$ denotes the ball of radius $r$ centered at $S$.
Proof of Prop. B.7.
Because the minimization problem does not depend on a quantity that is treated as random in Thm. 5.1, we may directly apply Thm. 5.1 to $k^{\star}_{T}$ to conclude with probability at least $1-\delta$
$|R^{\chi}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T}}}(h)|\leq L\rho_{k^{\star}_{T}}+\Delta\|w^{\textup{{{NN}}},k^{\star}_{T}}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$ (66)
We split into cases.
Case 1: $L\leq 1$
Because $L\leq 1$ and by the minimality of $k^{\star}_{T}$,
$L\rho_{k^{\star}_{T}}+\Delta\|w^{\textup{{{NN}}},k^{\star}_{T}}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$ $\leq\rho_{k^{\star}_{T}}+\Delta\|w^{\textup{{{NN}}},k^{\star}_{T}}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$ (67)
$=\min_{k\in T}\rho_{k}+\Delta\|w^{\textup{{NN}},k}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}.$ (68)
Case 2: $L>1$
Because $L>1$ and the second term is non-negative,
$L\rho_{k^{\star}_{T}}+\Delta\|w^{\textup{{{NN}}},k^{\star}_{T}}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$ $\leq L\left(\rho_{k^{\star}_{T}}+\Delta\|w^{\textup{{{NN}}},k^{\star}_{T}}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}\right)$ (69)
$=L\min_{k\in T}\rho_{k}+\Delta\|w^{\textup{{NN}},k}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}.$ (70)
Combining Eqns. 68, 70 and 66 gives the result. ∎
Proposition B.8 (Minimization over Powers of Two).
Let $T=\{1,\dots,N^{\textup{{val}}}\}$ and $T_{2}=\{2^{i}\}_{i=1}^{\lfloor\log_{2}N^{\textup{{val}}}\rfloor}$. Fix $\delta\in(0,1)$. Define $k^{\star}_{T_{2}}\in\arg\min_{k\in T_{2}}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}$ with $\beta_{\delta}=\Delta\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$. Define $C_{L}=\max(1,L)$. Under Assumptions 2.2, 2.1 and 2.5 with probability at least $1-\delta$
||| $R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq\sqrt{2}C_{L}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\right),$ (71)
where $B(S,r)$ denotes the ball of radius $r$ centered at $S$.
Proof.
Let $k^{\star}_{T}$ denote a minimizer of the bound on the right hand side, which exists since the minimization is over a finite set. If $k^{\star}_{T}=1$, we are done since $1\in T_{2}$. Otherwise, there exists a $\tilde{k}\in T_{2}$ such that,
$k^{\star}_{T}/2\leq\tilde{k}\leq k^{\star}_{T}.$ (72)
By monotonicity of the $k$th order fill distance in $k$,
$\rho_{\tilde{k}}\leq\rho_{k^{\star}_{T}}.$ (73)
This also implies,
$Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{\tilde{k}}))\leq Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k^{\star}_{T}})),$ (74)
since the measure of a subset is never larger than the measure of a set that contains it. Therefore,
$\min_{k\in T_{2}}\left(\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\right)$ $\leq\rho_{\tilde{k}}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{\tilde{k}})))}{\tilde{k}}}$ (75)
$\leq\rho_{k^{\star}_{T}}+\beta_{\delta}\sqrt{2}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k^{\star}_{T}}))}{k^{\star}_{T}}}$ (76)
$\leq\sqrt{2}(\rho_{k^{\star}_{T}}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k^{\star}_{T}}))}{k^{\star}_{T}}})$ (77)
$=\sqrt{2}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\right).$ (78)
The result now follows from Prop. B.7. ∎
B.5 Consistency of our Nearest Neighbor Method under Infill Asymptotics
In this section we prove Cor. 5.2, which establishes the consistency of SNN under infill asymptotics. The idea of the proof is to upper bound the $k$th order fill distance of the validation points in the test points to the fill distance of the validation points in $[0,1]^{d}$. This allows together with Thm. 5.1 and Prop. B.8 allows us to derive an upper bound on the error of our method in terms of the fill distance, from which the result follows.
B.5.1 Preliminary Lemma: Relating Fill Distances
Proposition B.9.
Let $A$ be an $\epsilon$-net for $[0,1]^{d}$. Then for $k\leq\left(\frac{1}{2\epsilon}\right)^{d}$
$\zeta^{k}(A,[0,1]^{d})\leq 2k^{1/d}\epsilon+\epsilon.$ (79)
Proof.
Let $S\in[0,1]^{d}$ and $\tau\in(0,1]$. Then,
$|A\cap B(S,\tau+\epsilon)|\geq\mathfrak{N}(\epsilon,B(S,\tau))$ (80)
because for any $S^{\prime}\in B(S,\tau)\cap[0,1]^{d}$, there is a point $a\in A$ such that $d(S^{\prime},a)\leq\epsilon$ and for any such point, this $a$ must also be in $B(s,\tau+\epsilon)$ by the triangle inequality. Let $C$ be an $\epsilon$-net of $B(s,\tau)\cap[0,1]^{d}$. Then by the definition of a net and by subadditivity
$\text{vol}(B(S,\tau)\cap[0,1]^{d})\leq\text{vol}(\cup_{c\in C}B(c,\epsilon))\leq|C|B_{d}\frac{1}{\epsilon^{d}}$ (81)
Since $S\in[0,1]^{d}$ and $\tau\leq 1$, at least one orthant of $B(s,\tau)$ is contained in $[0,1]^{d}$, so
$\text{vol}(B(S,\tau)\cap[0,1]^{d})\geq 2^{-d}B_{d}\frac{1}{\tau^{d}}.$ (82)
Combining the previous estimates,
$\mathfrak{N}(\epsilon,B(S,\tau)\cap[0,1]^{d})\geq\left(\frac{\tau}{2\epsilon}\right)^{d}$ (83)
Choose $\tau=2\epsilon k^{1/d}\in(0,1)$. Then
$|A\cap B(S,\tau+\epsilon)|\geq k.$ (84)
As $S$ was arbitrary, this holds for all $S\in[0,1]^{d}$, and so for all $S\in[0,1]^{d}$, there are $k$ points in $A$ in $B(S,2\epsilon k^{1/d}+\epsilon)$. ∎
B.5.2 Bound on Loss Depending on Fill Distance
We now present and prove an upper bound on the error of SNN that depends on the fill distance of the validation set in $[0,1]^{d}$. We will derive consistency of the estimator under infill asymptotics as a corollary of this bound. The bound will also be relevant in later discussion of model selection, where we are also interested in rates of convergence of estimators.
Corollary B.10 (Bound for Dense Validation Data and General Test Data).
Suppose that $\mathcal{S}=[0,1]^{d}$ and Assumptions 2.2, 2.1 and 2.5. Let $k^{\star}_{T_{2}}\in\arg\min_{k\in T_{2}}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}$ with $\beta_{\delta}=\Delta\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$. Then there exists a constant $K_{d,\delta,\Delta,L}$ such that with probability at least $1-\delta$
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq K_{d,\delta,\Delta,L}\tilde{\rho}^{\frac{d}{d+2}},$ (85)
where $\tilde{\rho}=\zeta(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},[0,1]^{d})$.
Proof.
For all, $\tilde{\rho}\geq 1$, the stated bound holds with $K=L\sqrt{d}+\beta_{\delta}$. Therefore, moving forward, we assume $\tilde{\rho}<1$.
Choose $k=\min(\lceil(\gamma\tilde{\rho})^{-\frac{2d}{d+2}}\rceil,N^{\textup{{val}}})$ for some $\gamma\in(0,1/2)$ to be specified later. Because $\frac{2d}{d+2}<d$ and $\gamma\leq\frac{1}{2}$, this $k$ satisfies the conditions of Prop. B.9 so,
$\rho_{k}\leq\zeta^{k}(A,[0,1]^{d})$ $\leq 2\min(\lceil(\gamma\tilde{\rho})^{-\frac{2d}{d+2}}\rceil^{1/d},(N^{\textup{{val}}})^{1/d})\tilde{\rho}+\tilde{\rho}$ (86)
$\leq 2(1+\gamma\tilde{\rho})^{-\frac{2}{d+2}})\tilde{\rho}+\tilde{\rho}$ (87)
$\leq 4\gamma^{-2}\tilde{\rho}^{\frac{d}{d+2}}.$ (88)
We now apply Thm. 5.1 and Prop. B.8 together with this bound to conclude with probability $1-\delta$
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|$ $\leq\sqrt{2}C_{L}(4\gamma^{-2}\tilde{\rho}^{\frac{d}{d+2}}+\frac{1}{\sqrt{k}}\beta_{\delta})$ (89)
$\leq\sqrt{2}C_{L}\left(4\gamma^{-2}\tilde{\rho}^{\frac{d}{d+2}}+\min\left(\gamma^{\frac{d}{d+2}}\tilde{\rho}^{\frac{d}{d+2}},\frac{1}{\sqrt{N^{\textup{{val}}}}}\right)\beta_{\delta}\right).$ (90)
Choosing $\gamma=\frac{1}{4}$ (for example) completes the proof. ∎
B.5.3 Consistency of Spatial Nearest Neighbors
We now restate and prove that the spatial nearest neighbor procedure we describe is consistent under infill asymptotics. This follows as a corollary of Cor. B.10 since the infill assumption means that for any fixed $\delta$, the upper bound in Cor. B.10 tends to zero with the fill distance.
See 5.2
Proof.
Take $\delta=\min(1,\tilde{\rho})$. Under infill asymptotics, this tends to zero. On the other hand, by Cor. B.10 (or more precisely Eqn. 90 which makes the depend of the bound on $\delta$ explicit), with probability at least $1-\delta$
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq K_{d,\Delta,L}\tilde{\rho}^{d/(d+2)}\sqrt{\log\frac{1}{\tilde{\rho}}}.$ (91)
The left hand side of this bound tends to zero with $\tilde{\rho}$, and so $\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)$ converges in probability to $R^{\chi}_{Q^{\textup{{test}}}}(h)$. ∎
B.6 Model Selection: Rates of Convergence of Spatial Nearest Neighbors
We now present an extended version of earlier discussion in Section 5.3, as well as results on rates of convergence of spatial nearest neighbors that provides some support to claims regarding model selection. As a special case, we consider grid prediction (for example for assessing the global performance of a map constructed using a predictive method). We then discuss rates of convergence in the more general setting earlier addressed in Cor. B.10.
B.6.1 A Heuristic Discussion of Model Selection with Increasing Amounts of Data
Suppose we have a fixed test task, and two data-driven algorithms for making predictions. We also suppose we have allocated a fixed percentage of data for training, and the remainder for validation. Can we use spatial nearest neighbors to select between the two method? As the amount of training data increases we would hope that both data-driven algorithms produce better predictive methods. We therefore do not expect consistency to be sufficient to select between the two sequences of predictive methods: if both are converging to the optimal estimator at (possibly different) rates as the amount of training data increases, we want our error in estimating the risk of the two methods to converge to zero faster at a rate faster in the amount of validation data than the rate that the slower converging method converges to the optimal predictor. This would suggest we should be able to reliably identify the better sequence of predictive methods as the amount of data increases.
We will assume we are in the additive, homoskedastic error setting so that $Y=f(S,\chi(S))+\epsilon$. We will assume that the noise is bounded, and that squared loss is used. Finally, we will assume that training data is also generated following this process. We will also assume that $f$ is Lipschitz continuous. In this setting, minimax pointwise regression rates are $\theta((N^{\textup{{train}}})^{-\frac{1}{d+2}})$ (cf. Tsybakov (2008, Theorem 2.3, Corollary 2.2) in one-dimension under a fixed grid design and Tibshirani (2023, Examples 3.1, 3.2) for a multi-dimensional version with both fixed grid and random design. The latter assumes Gaussian noise instead of bounded noise). Ideally, we would like a method for performing model selection to be able to distinguish between a sequence of predictive methods converging at slower than the minimax optimal rate and a sequence of predictive methods converging at the minimax rate and to be able to reliably select between two sequences of predictive methods both of which are converging at the minimax rate, but with different constants.
We will now present some finite sample and asymptotic bounds on the convergence of spatial nearest neighbors for a fixed hypothesis then return to the question of model selection in light of these results.
B.6.2 Rates of Convergence for Grid Prediction
We first consider the grid prediction task specifically, as this is a common problem in spatial analyses. For example, one might want to reconstruct air temperature across the continental United States on a dense grid (map) based on remotely sensed covariates observed on this grid and sparsely observed weather station data. We consider a test task grid prediction if the data falls on a regular $d$-dimensional grid.
Assumption B.11 (Grid prediction).
We say that a task is grid prediction if $\mathcal{S}=[0,1]^{d}$ and $Q^{\textup{{test}}}=\frac{1}{g^{d}}\sum_{S\in\{i/g:1\leq i\leq g\}^{d}}\delta_{S}$ for some $g\in\mathbb{N}$.
As long as the resolution of the map is high, both $1$-nearest neighbor and spatial nearest neighbors provide reliable estimates of the error.
Corollary B.12 (Bound on Estimation Error for Grid Prediction).
With the same assumptions as Thm. 5.1 and additionally Assumption B.11, with probability at least $1-\delta$
$\!\!\!|R^{\chi}_{Q^{\textup{{test}}}}(h)\!-\!\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}}(h)|\!\leq\!L\rho\!+\!\beta_{\delta}\sqrt{\max(\tfrac{2^{d}}{M^{\textup{{test}}}},(8\rho)^{d})}.$ (92)
Also, with probability at least $1-\delta$
$\!\!|R^{\chi}_{Q^{\textup{{test}}}}(h)\!-\!\!\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq\!C_{L}\!\!\left(\!\!\rho\!+\beta_{\delta}\sqrt{\max(\tfrac{2^{d}}{M^{\textup{{test}}}},(8\rho)^{d})}\right).$
with $\rho=\rho_{1},\beta_{\delta}=\Delta\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$ and $C_{L}=\max(1,L)$.
See Section B.6.5 for a proof. The right-hand side of the bound is small as long as there is a validation point near each test point and the resolution of the map is high. If the available data for validation is generated i.i.d. as in Prop. 3.3, then right hand side becomes, up to logarithmic factors in $N^{\textup{{val}}}$, $\frac{1}{\sqrt{M^{\textup{{test}}}}}+(N^{\textup{{val}}})^{\min(-1/2,-1/d)}$
B.6.3 Statement and Discussion of Result for IID validation data and Grid Prediction
Corollary B.13 (Convergence of Grid Prediction with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of Cor. B.12. Fix $\delta\in(0,1)$ and $k\in\{1,k^{\star}_{T_{2}}\}$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
$|\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)-R^{\chi}_{Q^{\textup{{test}}}}(h)|\leq K_{d,\delta,L,\Delta,c}\left((\tfrac{\log N^{\textup{{val}}}}{N^{\textup{{val}}}})^{\min(\frac{1}{2},\frac{1}{d})}+\tfrac{1}{\sqrt{M^{\textup{{test}}}}}\right).$
Proof.
From Prop. 3.3, there exists a constant $\gamma$ depending on $d,L,\delta,c,\Delta$ such that with probability at least $21-\delta/$ 2
$\rho\leq\gamma\left(\frac{\log N^{\textup{{val}}}}{N^{\textup{{val}}}}\right)^{1/d}.$ (93)
An upper bound on the bound in Cor. B.12 shows that for $k\in\{1,k^{\star}_{T_{2}}\}$ with probability $21-\delta/$ 2,
$|\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)-R^{\chi}_{Q^{\textup{{test}}}}(h)|\leq\gamma C_{L}\max(\beta_{\delta/2}8^{d/2},1)\left(\rho+\frac{1}{\sqrt{M^{\textup{{test}}}}}+\rho^{d/2}\right).$ (94)
Combining Eqn. 93 and Eqn. 94 via a union bound and using that $a+b\leq 2\max(a,b)$ completes the proof. ∎
This matches the bound proven in Portier et al. (2023, Proposition 3) which assumed that the test data was independent and identically distributed instead of on a regular grid. This rate of convergence is reasonably fast, particularly in low-dimensions. In particular ignoring the dependence of the bound on $M^{\textup{{test}}}$, which is reasonable as the number of test points in map prediction is often far larger than the number of available points for training and validation, it is faster than the minimax optimal rate of convergence for Lipschitz functions of $\theta((N^{\textup{{train}}})^{-1/(d+2)})$ discussed earlier. We therefore would expect both spatial nearest neighbors and $1$-nearest neighbors to perform well for model selection for grid prediction tasks if a fixed percentage of the data is used for training, and the remainder for validation. We emphasize we do not give a formal proof of this, just a heuristic argument suggesting why this should be the case. To give a formal proof would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
B.6.4 General Prediction Tasks
For general $Q^{\textup{{test}}}$ we can combine Cor. B.10 together with Prop. 3.3 to get some sense of the rate of convergence of the spatial nearest neighbor method if the validation data is independent and identically distributed form a measure with density lower bounded on $[0,1]^{d}$ and the test task is fixed.
Corollary B.14 (Convergence of Spatial Nearest Neighbor with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of Cor. B.10. Fix $\delta\in(0,1)$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
$|\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq K_{d,\delta,L,\Delta,c}\left((\tfrac{\log N^{\textup{{val}}}}{N^{\textup{{val}}}})\right)^{\frac{1}{d+2}}.$ (95)
Proof.
From Prop. 3.3, there exists a constant $\gamma$ depending on $d,L,\delta,c,\Delta$ such that with probability at least $21-\delta/$ 2
$\tilde{\rho}\leq\gamma\left(\frac{\log N^{\textup{{val}}}}{N^{\textup{{val}}}}\right)^{1/d}.$ (96)
Cor. B.10 implies that there exists a constant $K\geq 0$ such that with probability at least $21-\delta/$ 2
$|\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq K\tilde{\rho}^{\frac{d}{d+2}}.$ (97)
Combining Eqn. 96 and Eqn. 97 completes the proof. ∎
In this case, again up to logarithmic factors, Cor. B.14 means that SNN converges at the optimal rate of convergence for Lipschitz functions. This means we do not necessarily expect to be able to distinguish between two sequences of predictive methods that converge at the minimax rate, but we might expect to distinguish between two sequences of predictive methods if one converges to the optimal predictor at much slower than the minimax rate. We again emphasize that we do not formally show this, and to do so would would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
In contrast, both the holdout and $1$-nearest neighbor methods are not even always consistent for risk estimation in this setting, and therefore cannot be expected to reliably perform model selection.
B.6.5 Grid Prediction Proofs
In order to prove the claimed upper bound on grid prediction Cor. B.12 we use Thm. 5.1 (in particular Eqn. 13) together with an upper bound on the number of test points that lie within a ball of radius equal to the fill distance around any validation point. In order to do this, we will use that all the points in a grid are well-separated. We therefore begin by recalling the definition of a packing of a set, as well as a relationship between covering number and packing number.
Definition B.15 (Packing, Packing Number).
Let $A\subset\mathbb{R}^{d}$ a compact set. A (finite) set $B\subset A$ is called an $\epsilon$-packing of $A$ if for all $b,b^{\prime}\in B$, $\|b-b^{\prime}\|>\epsilon$. The $\epsilon$-packing number of a set $A$, $\mathfrak{M}(\epsilon,A)$ is the largest cardinality of an $\epsilon$-packing of $A$.
Proposition B.16 (Packing and Covering Numbers Wainwright 2019, Lemma 5.5).
For any $A\subset\mathbb{R}^{d}$ and $\epsilon>0$,
$\mathfrak{M}(2\epsilon,A)\leq\mathfrak{N}(\epsilon,A)\leq\mathfrak{M}(\epsilon,A).$ (98)
We now restate and prove Cor. B.12. See B.12
Proof.
The second inequality follows from the first by Prop. B.7 and because $1\in T_{2}$. We therefore focus on proving the case $k=1$.
In light of Thm. 5.1, it suffices to show that for these test location,
$M^{\textup{{test}}}Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho))=|B(S^{\textup{{val}}}_{n},\rho)\cap\{a/g:1\leq a\leq g\}^{d}|\leq\max\left(2^{d},8^{d}\rho^{d}M^{\textup{{test}}}\right).$ (99)
The set $\{a/g:1\leq a\leq g\}^{d}$ is a $\frac{1}{g+\epsilon}$-packing for any $\epsilon>0$, and so by the first inequality in Prop. B.16 and Prop. B.3
$|B(S^{\textup{{val}}}_{n},\rho)\cap\{a/g:1\leq a\leq g\}|\leq\mathfrak{M}(\tfrac{1}{(g+\epsilon)},B(S^{\textup{{val}}}_{n},\rho))\leq\mathfrak{N}(\tfrac{1}{2\rho(g+\epsilon)},B(0,1)).$ (100)
Applying Wainwright (2019, Lemma 5.7, Equation 5.9) and taking the limit as $\epsilon\to 0^{+}$
$\lim_{\epsilon\to 0^{+}}\mathfrak{N}(\tfrac{1}{2(g+\epsilon)\rho},B(0,1))\leq\lim_{\epsilon\to 0^{+}}(1+4(g+\epsilon)\rho)^{d}=(1+4g\rho)^{d}.$ (101)
By the binomial theorem and bounding the sum by the number of terms times the largest term
$(1+4g\rho)^{d}\leq 2^{d}\max(1,4^{d}g^{d}\rho^{d})=\max(2^{d},8^{d}M^{\textup{{test}}}\rho^{d}).$ (102)
Appendix C Additional Experimental Details
In this section, we provide additional details about the data, fitting procedures and validation procedures used in Section 6. Code used in experiments is available anonymously at: https://github.com/DavidRBurt/Consistent-Spatial-Validation. All code is implemented in Python3 (Van Rossum & Drake, 2009). Numpy is also heavily used for data generation and array manipulation (Harris et al., 2020).
C.1 Risk Estimation on Synthetic Data
In this section, we provide additional details and figures for our risk estimation experiment on synthetic data presented in Section 6.1. Section C.1.1 describes the process by which we generate both datasets considered and shows an example of the covariate and response spatial fields for each problem (Figs. 6 and 8). Section C.1.2 describes the procedure used to fit the predictive methods to each dataset. Section C.1.3 describes the implementation of the risk estimation procedures we compare. Section C.1.4 describes the metric reported, and Figs. 7 and 9 show the (signed) relative error of each estimator as we vary the amount of validation data available.
C.1.1 Simulation Data Generating Process Details
For both tasks, 100 datasets are generated following the process outlined below.
Grid data
Generation of training and validation sites: The first training point is selected via generating a point uniformly in $[-0.5,0.5]^{2}$, and making this the mean of a Gaussian mixture component, with standard deviation randomly sampled between 0.05 and 0.15. This mixture is initially given weight $1$, the first training point is then sampled from a Gaussian with this mean and standard deviation, the and weight of this mixture is increased to $2$. Subsequent points are sampled sequentially. For each $i$ between $2$ and the total number of training and validation points, a weight of $1$ is assigned to adding a mixture component. The new point is then sampled from either one of the existing mixture components, or the new mixture component, with probability proportional to the current weights assigned to each mixture component. The weight of the mixture from which the points, $w^{(t)}_{i(t)}$, is then increased as
$w^{(t+1)}_{i(t)}=w^{(t)}_{i(t)}+\frac{1}{w^{(t)}_{i(t)}}.$ (103)
This is reminiscent of a Chinese restaurant process (Pitman, 2006, Section 3.1), but the weights are increased more slowly, leading to more clusters being formed and less large clusters typically.
If a new mixture component is generated, a mean for the mixture component is generated on $[-0.5,0.5]^{2}$, and a standard deviation is selected uniformly on $[0.05,0.15]$. Conditional on the mixture component, the new point is sampled from a Gaussian distribution with the components mean and standard deviation.
The first $1000$ points generated this way are taken to be the training data, and the remaining $N^{\textup{{val}}}$ points generated this way are the validation data. An example of the training and validation data generated through this process are shown in the top left of Fig. 6.
Generation of test data The test data is $\{(-0.5+a/49,-0.5+b/49):0\leq a,b\leq 49\}$. That is, it is a regular grid on $[-0.5,0.5]^{2}$.
Generation of Covariates
The covariates are generated as a zero-mean Gaussian process with an isotropic Matérn 3/2 covariance function with lengthscale $0.3$ and scale parameter $1$. That is, the covariance function is,
$k_{\chi}(S,S^{\prime})=\left(1+\frac{\sqrt{3}\|S-S^{\prime}\|_{2}}{0.3}\right)\exp\left(-\frac{\sqrt{3}\|S-S^{\prime}\|_{2}}{0.3}\right).$ (104)
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using tensorflow probability (Dillon et al., 2017). We generate two covariates spatial processes via this process $\chi=(\chi^{(1)},\chi^{(2)})$.
Generation of Response Once the sites and covariates have been generated, the response variable is sampled from a Gaussian process with zero mean. The covariance function of the Gaussian is a sum of two, $2$ dimensional isotropic Matérn 3/2 kernels:
$\!\!k(S,S^{\prime})\!=0.5\left(1+\tfrac{\sqrt{3}\|S-S^{\prime}\|_{2}}{0.5}\right)\exp\!\left(-\tfrac{\sqrt{3}\|S-S^{\prime}\|_{2}}{0.5}\right)\!+\!\left(1\!+\sqrt{3}\|\chi(S)-\chi(S^{\prime})\|_{2}\right)\exp\!\left(-\sqrt{3}\|\chi(S)-\chi(S^{\prime})\|_{2}\right).$ (105)
Independent, identically distributed Gaussian noise is added to the function values with variance $0.1$.
Figure 6: Data for the grid prediction task. An example of a single sample of the training and validation points (with 1000 training points and 250 validation points) is shown in the top left. The top right and bottom left show the covariates as a function of space, while the bottom right shows the mean of the response variable as a function of space.
Point Prediction Task
Generation of training and validation sites: The training and validation points are sampled independently and identically from a uniform distribution supported on $[-0.5,0.5]^{2}$. $1000$ training points are used in all experiments. The number of validation points is varied in $\{250\times 2^{\ell}\}_{\ell=0}^{5}$.
Generation of test site The test site is fixed to be the origin. We generate 2500 response values at the origin so that when we compute the empirical risk we expect it to accurately reflect that actual risk.
Generation of covariates: The covariates are generated as a zero-mean Gaussian process with an isotropic squared exponential covariance function with lengthscale $0.3$ and scale parameter $1$. That is, the covariance function is,
$k_{\chi}(S,S^{\prime})=\exp\left(-\frac{\|S-S^{\prime}\|_{2}^{2}}{2\cdot 0.3^{2}}\right).$ (106)
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using tensorflow probability (Dillon et al., 2017). We generate two covariates via this process $X=(X^{(1)},X^{(2)})$. Generation of response Once the sites and covariates have been generated, the response variable is sampled from a Gaussian process with zero mean. The covariance function of the Gaussian is a sum of two, $2$ dimensional isotropic squared exponential kernels:
$k(S,S^{\prime})=0.5\exp\left(-\frac{\|S-S^{\prime}\|^{2}_{2}}{2\cdot 0.5^{2}}\right)+\exp\left(-\frac{\|\chi(S)-\chi(S^{\prime})\|_{2}}{2}\right).$ (107)
Independent, identically distributed Gaussian noise is added to the function values with variance $0.1$.
C.1.2 Model Fitting
We fit a Gaussian process regression model to using only the first covariate $\chi^{(1)}$ to the training data. The prior is taken to be the same as the data generating process, but with (only) $\chi^{(1)}$ in place of $(\chi^{(1)},\chi^{(2)})$ in the second kernel in Eqn. 105 and Eqn. 107 for the two datasets respectively. The mean of the posterior process is used for predictions, and is calculated using GPFlow (Matthews et al., 2017).
C.1.3 Implementation of Risk Estimation
The holdout is implemented by taking an (unweighted) average of the loss on each validation point. Both nearest neighbor methods are implemented using scikit-learn (Pedregosa et al., 2011) with $kd$-trees and Euclidean distance. For $k^{\star}_{T_{2}}$ nearest neighbors, nearest neighbors is performed for all $k$ that are powers of $2$ less than $N^{\textup{{val}}}$, and the value of $k$ with the smallest bound is used for risk estimation. This is done with $\delta=0.1$ and $\Delta=1$. Table 2 and Table 3 show the values of $k$ chosen for grid and point prediction respectively. For the grid prediction task, $k^{\star}_{T_{2}}$ tends to be small, even as the size of the validation set becomes large. This supported by Cor. B.12, since even $1$-nearest neighbor reliably estimates risk in this setting. For the point prediction task, $k^{\star}_{T_{2}}$ grows with $N^{\textup{{val}}}$.
Table 2: Value of $k^{\star}_{T_{2}}$ chosen by minimizing the bound in the grid prediction task. Because the test data is well separated, the variance of the estimator is small even when $k$ is small. As a result the value of $k$ that minimizes the upper bound is generally small, even as the number of validation points increases.
Number of Validation Points
$\mathbf{k^{\star}_{T_{2}}}$ 250 500 1000 2000 4000 8000
1 78 83 79 91 90 80
2 19 16 21 9 10 20
4 3 1 0 0 0 0
Table 3: Value of $k^{\star}_{T_{2}}$ chosen by minimizing the bound in the point prediction task. In this task, there is generally a bias-variance trade-off that must be balances. As a result the value of $k$ that minimizes the upper bound increases as the amount of available validation data increases..
Number of Validation Points
$\mathbf{k^{\star}_{T_{2}}}$ 250 500 1000 2000 4000 8000
16 1 0 0 0 0 0
32 97 27 0 0 0 0
64 2 73 98 24 0 0
128 0 0 2 76 100 11
256 0 0 0 0 0 89
C.1.4 Metrics Reported and Additional Figures
We use truncated, squared loss,
$\ell(a,b)=\min(1.0,(a-b)^{2}),$ (108)
which is bounded by $1.0$. The empirical risk is calculated as in Eqn. 15. In both cases, we have $2500$ test points that are averaged over; as this is a simple Monte Carlo estimate, of a bounded random variable, we expect this to be concentrated around the mean. Formally, we may apply Hoeffding’s inequality to conclude that with probability at least $0.95$
$|\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)-R^{\chi}_{Q^{\textup{{test}}}}(h)|\leq 1.0\sqrt{\frac{1}{2\times 2500}\log\frac{2}{0.05}}\approx 0.0272.$ (109)
Figs. 7 and 9 show the relative errors of each estimation, calculated as
$\frac{\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}(h)}{\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)}.$ (110)
From this, we can see that the holdout method has a bias in both cases that does not appear to go away as the number of validation points increases. In contrast, the $1$-nearest neighbor method primarily suffers due to a variance issue when it fails to converge. We again see in both instances the $k^{\star}_{T_{2}}$-nearest neighbor approach appears to concentrate around zero error as the number of validation points increases.
Figure 7: The relative error in estimating the empirical risk for each method is plotted against the number of validation points used for the grid prediction task. The holdout is biased, even for large $N^{\textup{{val}}}$. The $1$-nearest neighbor and $k^{\star}_{T_{2}}$-nearest neighbor estimates both have small relative error for large $N^{\textup{{val}}}$.
Figure 8: Data for the point prediction task. An example of a single sample of the training and validation points (with 1000 training points and 250 validation points) is shown in the top left. The top right and bottom left show the covariates as a function of space, while the bottom right shows the mean of the response variable as a function of space.
Figure 9: The relative error in estimating the empirical risk for each method is plotted against the number of validation points used for the point prediction task. The holdout has a small but non-negligible bias. The $1$-nearest neighbor method has a small bias, but large variance. The nearest neighbor procedure with $k^{\star}_{T_{2}}$ neighbors has small relative error for large $N^{\textup{{val}}}$.
C.2 Details for Model Selection Experiment
We present additional details of the experiment discusses in Section 6.2.
Data Generation
The data generation is fully described by Eqn. 16.
Model Fitting
We consider two models. The first is a constant predictor that predicts $0.25$. The second is an affine model (a linear model with an intercept) fit by minimizing the mean absolute error from the line to the training points. This is fit using the Scikit-learn quantile regression with the (default) “HiGHS” solver (Huangfu & Hall, 2015).
Estimation of Risk
The validation estimates used are calculated in the same as the synthetic experiments outlined previously in Section C.1.3. We use $\Delta=1$ in the bound when selecting $k^{\star}_{T_{2}}$, even though the absolute value loss used can be larger than $1$. We don’t expect this to have a significant impact on the results, as the upper bound we minimize is already misspecified in a similar way by not using the actual Lipschitz constant of the function. We again use $\delta=0.1$ when selecting $k^{\star}_{T_{2}}$.
C.3 Details of Air Temperature Prediction Experiment
We now present additional details of the air temperature experiment presented in Section 6.3.
C.3.1 Results for Grid Prediction
Table 4 shows the results for holdout, $1$-nearest neighbor and SNN with the test set each grid point in the map that is located in the continental United States. We see good agreement between all three method. This is expect for $1$-nearest neighbor and SNN.
Table 4: Estimates of risk given by each method. All three methods agree reasonably well (within ±$\pm 2$ standard deviation of the estimate given by holdout for both geographically weighted regression and spatial regression in this task. In particular, all three methods suggest that the geographically weighted regression method has lower risk on this task.
Holdout 1NN SNN
GWR $\mathbf{0.84\pm 0.03}$ $\mathbf{0.83}$ $\mathbf{0.83}$
Spatial GP $0.96\pm 0.04$ $0.92$ $0.92$
C.3.2 Data Sources
The land surface temperature used is from MODIS Aqua (Wan et al., 2021) and is monthly average land surface temperature on a $0.05$ degree grid. We download monthly average weather station data from the Global Historical Climatology Network (Menne et al., 2018). Latitude and longitude of major United States (US) urban areas are from the 2023 US census gazetteer (United States Census Bureau, 2023). All of these datasets are produced by US government agencies and are openly available for use.
C.3.3 Data Pre-processing
We assign the land surface temperature at the nearest point (using a spherical approximation to distance between points) to each weather station. The nearest point is found using the scikit-learn implementation of nearest neighbor algorithm using the ‘ball-tree’ (Omohundro, 2009) and ‘Haversine’ metric. Temperatures are converted to Celsius from Kelvin. We remove all rows where the Land Surface temperature is not available. We use the weather station data uploaded to GHCNM as of January 15, 2024. We filter out weather stations outside of the United States (based on the station ID). We also filter out stations with a non-empty quality control flag or no temperature recorded for January 2018 (the month we consider). Finally, we remove stations in Hawaii or Alaska to restrict the station to the continental United States. In total there are 6422 weather stations. We use 70% of the stations for fitting the models, and holdout the remaining 30% of stations for calculating each of the validation procedures. When building the test data, we remove points outside the United States based on a reverse geocoding lookup with Thampi (2015) to the nearest city. This does not create an exact boundary (since it is based on the nearest city or town and not the country in which the latitude and longitude is based in) but is a good proxy for whether or not a point is in the United States. Fig. 10 of the available weather stations for model fitting and validation, colored by monthly average temperature in January 2018.
Figure 10: Weather stations used in the air prediction temperature task we considered, colored by temperature in January 2018 in degrees Celsius.
C.3.4 Model Fitting
Inspired by Hooker et al. (2018), we fit a geographically weighted least squares regression model using the land surface temperature at day and night. In particular, we fit an affine model, with the coefficients, $\beta(S)$ depending on the location that will be predicted at. $\beta(S)$ is selected by minimizing solving the weighted least squares problem,
$\hat{\beta}(S)\in\arg\min_{(b_{0},b_{1},b_{2})\in\mathbb{R}^{3}}\sum_{i=1}^{n^{\text{train}}}w_{i}(S)(Y^{\text{train}}_{i}-(b_{0}+b_{1}X^{\text{train},1}_{i}+b_{2}X^{\text{train},2}_{i}))^{2},$ (111)
[…]
We also consider a simple baseline model fit using only the weather station data. We fit a Gaussian process with zero prior mean and Matérn 3/2 kernel to the weather stations with covariate the spatial locations in latitude and longitude converted to radians, and a Gaussian likelihood model. We fit the parameters of the kernel using L-BFGS to attempt to maximize the marginal likelihood of the parameters. The parameters fit are two lengthscale parameters (one for each spatial dimension), a kernel scale parameter, and a likelihood variance parameter. The mean is removed from the training data prior to fitting, the kernel lengthscales are set to to standard deviation of each covariate and the kernel variance parameter is set to equal the variance of the training response data, and the likelihood variance parameter is set to equal $0.1$-times the variance of the training response variable. A maximum of $15$ iterations of L-BFGS are run.
C.4 Risk Estimation Details
The holdout is implemented as in previous experiments (Section C.1.3). We estimate the standard error of the holdout empirically by computing the sample standard deviation of the sum of the losses,
$\hat{\sigma}^{2}=\left(\frac{1}{N^{\textup{{val}}}(N^{\textup{{val}}}-1)}\sum_{j=1}^{N^{\textup{{val}}}}\ell(Y^{\textup{{val}}}_{j},h^{\chi}(S^{\textup{{val}}}_{j}))^{2}\right)^{1/2}.$ (112)
Table 1 reports the holdout estimate $\pm$ two standard deviation.
The nearest neighbor methods are implemented using the scikit-learn implementation with Haversine distance and the ball-tree algorithm. $k^{\star}_{T_{2}}$ is selected with $\delta=0.1$ and $\Delta=1$ even though this is not an upper bound on the loss. The Haversine distance is used directly, not rescaled to kilometers.